### v1

# The Role of Task Complexity in Emergent Abilities of Small Language Models

Abstract
We investigate the relationship between task complexity and the minimum model size required for learning specific tasks in small transformer models. We focus on the ListOps dataset, consisting of nested math operations. We define the task complexity as the Kolmogorov complexity (KC) of the code solving the task, using a rough proxy for KC. We find a power-law relation between KC and parameters required to learn, suggesting number of parameters to learn harder task increases almost cubic in KC. On individual math operations, sum mod 10 is hardest to learn. Surprisingly, when combining tasks, we observe that sum is learned earlier and with fewer parameters when trained alongside max and median. Analyzing the model, we find strong evidence that models trained on sum alone and models trained jointly converge to different algorithms. Concretely, the sum alone model doesn’t seem to have learned number properties in the embedding layer, likely memorizing the sum table. In contrast, models trained on three tasks (maximum, median and sum) reveals that joint training results in clear number-like properties. Finally, we also find evidence that the sum-only model utilizes its feedforward layer more than the jointly trained model. Conversely, the attention layer in the joint model is activated more than the sum model. Our findings suggest there is another dimension to emergent abilities in language models, namely the algorithms being learned, potentially impacting scaling laws.
[…]
1 Related works
Scaling laws for language models characterize how validation loss scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design. However, this perspective overlooks critical aspects of learning. Our results show that the composition of the training data—specifically, the mix of tasks—can significantly alter what and how a model learns. Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
In general, language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore, it allows us to tune the task difficulty by combining different operations.
Here, we analyze how language models learn arithmetic operations by training small models on the ListOps dataset. The ListOps dataset allows us to measure model accuracy explicitly by asking the model to solve mathematical equations. Using this bottom-up approach, we show that once the number of trainable parameters in the model reaches a critical point, the model is able to learn the task. When training models on datasets based on different levels of operation difficulty, this critical point shifts according to the complexity of the task. Contradictorily, we find that combining operations makes the task easier for the model than individual operations, suggesting that task diversity plays a significant role.
lubana2024percolation discuss that structure learning drives emergent abilities. Seems very relevant. Emergence of hidden capabilities park2024emergence Another grokking paper liu2022towards claims grokking is between “ comprehension” and “memorization”. We may have more evidence for that.
2 Introduction
Scaling laws for language models kaplan2020scaling characterize how performance scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design. Moreover, study of emergent abilities in LLM showed wei2022emergent that various capabilities emerge at different model sizes. Both of these directions suggest an intimate and universal relations between model size and the emergence of abilities. However, these views miss important details about the role of training curriculum.
In general language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore, it allows us to tune the task difficulty by combining different operations.
There is always more than one way to solve a task. Each solution might require dramatically different amount of compute and memory resources. If models learn to implement specific solution algorithms, variability in resource requirements should be reflected in some way in the model size or other metrics. Oftentimes, when humans try to solve a task, the simplest solution (meaning, easiest to find) is a brute-force one and not particularly efficient. Given more thought, knowledge and experience, we often manage to find more efficient solutions. So a key question to ask is whether the same type of algorithmic change can be observed in machine learning models and how it can be controlled.
How do we know which solution path a model took? We explore this problem through the lens of basic math operations, using the setting of the ListOps dataset (nangia2018listops). We show strong evidence for learning curriculum dramatically changing the algorithms learned by the models. Models of the same size may follow different learning paths depending on the training setup, suggesting that the tasks themselves influence the internal algorithms that emerge during training.
It is worth noting that our observations point to more than just task complexity being the driving force. Complexity of a task may be defined using Kolmogorov complexity (KC) kolmogorov1965three; li2008introduction, defined as the length of the shortest code producing a desired output. What we show is that models may require vastly different minimum sizes to learn the same task, depending on the training curriculum. Additionally, we show strong evidence of a clear distinction between the internal algorithm learned by the models using different curricula. Such a distinct algorithm could represent learning different algorithms to solve the same problem, with each algorithm having a different resource requirement, e.g. brute-force versus an efficient code. Hence we argue the model size depends more on the algorithmic complexity, or description length of a specific algorithm, rather than the KC of the task. This also raises an important question: what training regimes are most effective at inducing the learning of more efficient algorithms?
Using the ListOps dataset as our experimental framework, we investigate how small transformer models learn nested mathematical operations—specifically maximum (MAX), minimum (MIN), median (MED), and sum modulo $n$ (SUM). We train the models on a single or a mixture of these operations. We make the following observations:
  1. 1.
    A clear hierarchy in the model size vs learning transition: MAX and MIN are easiest to learn. Introducing MED and all combinations of MAX, MIN and MED require significantly more parameters to learn (2$\sim$ 2$\times$ 2 × for $n=26$). Next, adding SUM (mod $n$) to the mix makes it again significantly harder (1.5$\sim$ 1.5$\times$ 1.5 × for $n=26$).
  2. 2.
    SUM is the hardest: Paradoxically, the clear outlier as the hardest task to learn was SUM alone, not the more complex mixture of SUM and other operations.
  3. 3.
    Strong evidence for algorithmic diversity: The embedding layer of the pure SUM models, suggests models not learning the properties of numbers, whereas, mixed SUM models show strong number features including ordering and parity, with parity only observed in mixed SUM tasks.
  4. 4.
    Pretraining with MAX+MED makes SUM easier: Switching from MAX+MED to SUM has the same effect as mixed task, with much smaller models being able to learn SUM.
We hypothesize that compared to the pure SUM task, the mixed SUM training makes it easier for the models to discover number properties of the symbols. This seems to lead the mixed SUM models learning an entirely different algorithm than the pure SUM model. It may be that the pure SUM models are memorizing the table of additions. Indeed, we observe a similar difficulty to learning pure SUM for models learning a scrambled SUM dataset, where we create a scrambled sum table with the right hand side of $A+B=C$ being shuffled. In the scrambled SUM task, any number relation between the symbols has been erased and the easiest way to learn it should be to memorize the sum table, rather than overfit the random patterns in the sum table.
These results provide compelling evidence that joint training guides models towards finding alternative, more efficient solutions. They suggest there is more nuance to the scaling laws and they can be significantly affected by the training curriculum and strategies.
3 Methodology
We train small-scale transformer models and analyze both their performance and internal representations. This section outlines our methodology, detailing the dataset, model architecture, and evaluation protocols.
Choice of task
ListOps consists of nested mathematical equations involving operations. Our setup uses MAX, MIN, MED (median), and SUM (sum modulo $n$) applied to numbers (0-$n-1$). We conduct experiments using both $n=10$ and $n=26$, in which case we use capital letters A…Z to represent $0\dots 25$ (Appendix D.1). Choosing ListOps was motivated by several key factors that align with our research objectives:
  1. 1.
    Procedural Generation: ListOps is a synthetically generated dataset, allowing us to create a large volume of examples with controlled difficulty.
  2. 2.
    Exact Evaluation: The mathematical nature of ListOps operations ensures unambiguous, exact evaluation of model outputs.
  3. 3.
    Adjustable Task Complexity: ListOps offers a framework to modulate problem complexity by combining different operations and adjusting nesting depths.
  4. 4.
    Inter-task Relationships: The dataset’s multiple operations provide an opportunity to explore task synergies and interference.
Dataset Description
[…]
Tokenization and CoT:
We employ a character-based tokenization strategy for processing ListOps expressions. (Appendix D.2). We find that directly solving nested ListOps in one step can be quite challenging for transformer model (Fig. 22) Even with a maximum of three nesting levels with three operands (inputs) we find that GPT models with over 10 million parameters still fail to learn the task. To enhance model performance, particularly on more complex operations like sum modulo 10, we introduced a chain of thought (CoT) approach in our training data (Appendix D.3): s(%12(%34))>(%127)>0=0e, where the ‘>’ token means one step of CoT, wherein we solve the right-most, inner-most parenthesis. This is similar to the scratchpad used in lee2024teaching, among others. The tasks become increasingly challenging to learn with increasing nesting, number of operands, and length in a predictable manner (Appendix Fig. 27).
Train Test Split:
The CoT could result in many patterns appearing in one equation. To ensure that all the patterns in the test set are not also present in the training data, we select 100 triplets of numbers, e.g. $749$, to make an exclusion set. The training data is chosen from equations which never encounter the triplets in the exclusion set, and the test set is chosen such that each equation contains at least one excluded triplet.
Performance Evaluation:
We randomly select 1000 equations from the held-out test set. The model is prompted to generate solutions character-by-character, starting from the equation prompt (e.g., s(+3(-749))>). We use the output to produce two metrics: Loss: using cross-entropy computed for every character of the output. Accuracy: based on the number of correct answers evaluated using only the final answer, which we define as the first character after the first ‘$=$’ symbol.
Model Architecture
For our experiments, we employed a series of tiny GPT models, inspired by the nanoGPT architecture (karpathy2022nanogpt). Each model in our study uses a single attention head. We set the feedforward hidden dimension to four times the embedding dimension, a common practice in transformer architectures, providing sufficient complexity in the feedforward networks while keeping the model size constrained. By varying the embedding size and the number of transformer layers, we can create a range of models with different parameter counts.
Figure 1: Emergence of abilities in ListOps: Each plot shows the same group of small transformer models trained on a different mix of the four operations MAX, MIN, MED, and SUM. Red dots are models reaching more than 50% accuracy, and blue dots are less than 50%. The dashed green line is a logistic fit, and the yellow star indicates the transition point at 50%. The x-axis is the model size (number of parameters), and the plots are sorted in ascending order of transition points. The bottom panel shows a bar plot of the model sizes at the transition points, with each group distinguished by a different color.
4 Results
To understand how language models acquire mathematical abilities, we focus on accuracy as the primary metric for observing the emergence of learning, following the approach of the emergent abilities literature as in wei2022emergent. While some subsequent studies have questioned the concept of emergence by examining other metrics (schaeffer2024emergent), we argue that this critique overlooks similar patterns observed in physics: In a phase transition, only some quantities may exhibit discontinuous changes. A metric that can capture the emergence of the new phase is called the “order parameter”. Hence, we choose to use accuracy as our order parameter.
We run experiments with embedding dimension $n_{embed}$ between 4–128, and number of layers 1–4. In most settings, the single-layer transformers never reached 100% accuracy, so we decided to omit them. Similar to the emergence literature, we observe the total number of parameters to be the strongest indicator of accuracy, rather than $n_{embed}$ or depth. We mainly show results for modulo 26 experiments, as the patterns are more conclusive, but we also observed similar patterns for mod 10 (Appendix A.2).
Transition to Learning.
Figure 1 shows the transition to learning for all the 15 combinations of the four operations. The bottom plot shows the transition points, found using a logistic fit. The tasks involving MED are colored in green, and tasks involving SUM are colored in red, except for the pure SUM task, which is the curious outlier. We observe a clear hierarchy of tasks: adding MED to the mix makes the task significantly harder, while tasks containing SUM are even harder than MED.
However, surprisingly, the hardest task by far is the pure SUM, requiring at least twice as many parameters as even the mixed SUM tasks. This is peculiar, as one might guess the complexity of the task would depend on the number of tasks mixed together and the hardness of each component task. Again, counter this argument, for mixed SUM tasks we observe the exact opposite: the mixture of all four tasks is easier than mixture of three, which is easier than two. So, something must be happening when we train multiple math tasks jointly, which makes it easier for models to learn. Such behavior has also been reported in prior works on language models lee2024teaching where task diversity is observed to make models smarter. We want to get to the bottom of how this mixture is leading to a dramatic reduction in the learning transition when training on mixed tasks in ListOps.
Figure 2: PCA of embeddings: We choose all models which reached over $90$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure MAX, or all mixtures involving a given operation, e.g. Mixed SUM. Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do. Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
Embedding Layer.
To understand the difference between the hierarchy of transition points observed, we first examine the embedding layer of the models which learned the tasks (acc. $>90\%$) (see Appendices A.1 and A.2 for more). While the $n_{vocab}\times d_{embed}$ embedding layers of models of different sizes cannot be directly compared, we can compare them using $n_{vocab}$ dimensional principal component analysis (PCA). Moreover, while the PCs of individual models may be noisy, we can find aggregate PCs by first computing the $n_{vocab}\times n_{vocab}$ correlation matrices of a group of models and averaging them. This approach not only reduces the noise but also allows us to combine models of different sizes.
Figure 3: PC5 and 6 of Average Mixed SUM embedding correlations perfectly separate odd and even numbers.
An interesting picture emerges when we average models which learned each individual operation. Figure 2 shows the correlation matrix and top PCs for the embeddings of models trained on the groups we identified in the transition plot: 1) Pure MIN and MAX, 2) Pure and Mixed MED excluding mixture with SUM, 3) Mixed SUM, and 4) Pure SUM. We note that PCs for group 1 (pure MAX or MIN), as well as the mixture groups 2 and 3, exhibit clear, regular patterns. In contrast, the PCs for group 4 (pure SUM) look random.
We also observe that the regular patterns are not all the same. For group 1, 2 and 3, the top three PCs are almost identical, with PC1 always sorting numbers, and PC2 and PC3 being the two lowest frequency waves (1 and 2 peaks). These three are all top eigenvector of the Laplacian of a weighted line graph. Interestingly, PC4 and PC5 differ among these models. In groups 2 and 3, MED and Mixed SUM, PC4 seems close to the next frequency (3 peaks). For PC5, in group 1 it looks random, in 2 it looks like the next wave’s frequency (4 peaks), but for group 3, it does something striking: it identifies parity by separating odd from even numbers. More specifically, a linear combination of PC5 and PC6 does parity (Fig. 3). The presence of the parity mode in the embedding turns out to be a very robust feature of models trained on Mixed SUM tasks. We observe it not only in this grouping, but even very prominently in individual models trained on MAX+MED+SUM (Fig. 4) and mod 10 (Fig. 13, 20). Based on the observations, we pose the following hypotheses:
  1. 1.
    Sorting-based MAX, MIN, MED: PCA in these models show strong features of a line graph where consecutive numbers have edges and have an ordering.
  2. 2.
    Lookup-based Pure SUM: The lack of regularity in PCs suggest a lack of understanding of number properties, hence the algorithm might be lookup and memory-based.
  3. 3.
    Parity and sorting-based solution to Mixed SUM: Parity appears to play a very prominent role in the Mixed SUM algorithm learned.
Considering that much smaller models learned Mixed SUM compared to Pure SUM, we hypothesize that models in these two groups learned entirely different algorithms for SUM. Mixed training seems to lead the model to develop some understanding of number properties, allowing it to use efficient and compact algorithms to solve SUM. Pure SUM models, in contrast, seem to be treating numbers like random symbols. If this is the case, we believe the models need to memorize a sum table, storing $(A,B)\to C$ for $A+B=C$. One way to test the hypothesis of Pure SUM finding symbolic solutions is to design a similar problem where combination of symbols $(A,B)$ turns into $C$, analogous to a shuffled sum table.
Shuffled SUM.
We conduct experiments with a shuffled, symmetric SUM table (i.e. $A+B=B+A$) in Mod 10, one set with Pure SUM and one on MAX+MED+Shuffled SUM. We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version (Fig. 15). Additionally, we observe that MAX+MED+Shuffled SUM is again more difficult than all operations except pure sum, suggesting that the number properties played an important role in the other tasks becoming easier in mixed training. The PCA lacks most of the features we observed in regular numbers, save for partial ordering, as MAX and MED still use the normal number ordering (Fig. 16). Coloring the numbers based on parity, we don’t see the clear parity separation we observe in the regular mod 10 experiments (Fig. 11), as expected.
Figure 4: Evolution of training loss, accuracy, and PCs of the embedding layer, Mod 26: The top shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 96 and 3 layers, trained either on pure SUM (blue) or on mixed MAX+MED+SUM (red). Models trained on MAX+MED+SUM were evaluated separately on pure MAX, MED, and SUM subsets; the corresponding accuracy curves are shown in green, purple, and orange, respectively. Curves are the mean on 3 runs, with shaded ±$\pm\sigma$. All models were trained for 20000 iterations. The figures beneath the main plot display PCA embeddings revealing that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss. The also show a prominent parity separation emerging in PC2 and PC4. Parity is colored by red and blue. In contrast, models trained solely on SUM exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
Learning dynamics Pure SUM vs MAX+MED+SUM.
Another piece of evidence for the difference between SUM and the rest comes from the training dynamics. Figure 4 shows the loss curves and the evolution of test accuracy in a pure SUM model vs MAX+MED+SUM, for models with $n_{embed}=96$ and 3 layers, which is slightly above the pure SUM learning transition. The mixed model was evaluated separately for each of the MAX, MED, and SUM test sets to see when each ability emerged. We observed a subtle but important sequence in learning: MAX was learned fastest, followed closely by MED, and then SUM. Crucially, the gaps between the learning of these operations were minimal, with the model beginning to grasp SUM while still perfecting MED. The accuracy curves also show that joint training accelerates the learning of the SUM operation, as evidenced by the much earlier rise of the orange dashed line (joint training) compared to the blue dashed line (pure SUM training), occurring at almost half the number of steps of pure SUM.
Joint Training May Shrink the Search Space.
The model trained on mixed data can take a very different route and converge to a different solution. We see that training on MAX, MIN, or MED all lead to embeddings which exhibit number properties. Because of the CoT steps, the model’s loss decreases when it learns any of the operations involved in the mix. For instance, since learning MAX is much easier than SUM, gradient descent may first learn to solve MAX, which is strongly associated with learning a representation for the digits with the correct ordering. This restriction of the embedding layer can make it significantly easier for the model to learn other number properties and, possibly, find a number-based algorithm for SUM, exploiting the properties of numbers instead of memorizing symbolic patterns.
4.1 Testing the embedding restriction hypothesis.
Figure 5: Learning SUM by finetuning MAX+MED: We train model much smaller than the sum-only learning transition (48 embedding, 2 layers). By switching the training data slowly from MAX+MED to pure SUM (never showing expression mixing all three) the model is able to learn SUM (blue) in this much lower parameter regime. In comparison, the pure SUM models (red) did not learn at this size.
We test the hypothesis that restricting the embedding to what is learned by operations like MAX and MED could lead to smaller models learning SUM. To further explore this hypothesis, we designed a transfer learning experiment: 1) We first trained a model ($n_{embed}=48$ and three layers, which on pure SUM was unable to learn) on MAX and MED operations until proficiency. 2) We then gradually introduced SUM operations, increasing their proportion in the training data from 0% to 100% over 1000 steps, while simultaneously phasing out MAX and MED.
Key findings from this experiment include:
  • The model began learning SUM immediately upon its introduction, despite never seeing mixed expressions (e.g., SUM with MAX or MED).
  • Interestingly, the model started to forget MAX and MED once these operations were no longer present in the training data.
  • Crucially, we verified that the model retained its number-like embedding structure even after MAX and MED were completely phased out.
Perhaps most strikingly, we found that even much smaller models (embedding size 24) could learn SUM perfectly using this hybrid approach, mirroring the efficiency of MAX+MED+SUM models. This resulted in a model capable of performing SUM operations that was 7x smaller than the PURE SUM model, while relying on a more sophisticated, number-like internal representation.
Attention and feedforward layers.
To investigate the internal dynamics of the sum-only vs MAX+MED+SUM (all3) models, we focused on the final layer of a 3-layer transformer network, featuring a single attention head and an embedding dimension of 128 (Appendix E). We find:
  1. 1.
    In the attention sublayer, the MAX+MED+SUM model shows slightly higher ratio values, suggesting that attention mechanisms play a more pronounced role when the model is trained on diverse operations. This could indicate that the attention sublayer is capturing more complex patterns or relationships necessary for handling multiple operations.
  2. 2.
    Conversely, in the feedforward sublayer, the SUM model demonstrates significantly higher ratio values. This suggests that when trained on SUM alone, the model relies more heavily on the feedforward network for computation. This could imply that the SUM operation is being implemented more directly through feedforward transformations.
  3. 3.
    The larger effect size in the feedforward layer (-0.34) compared to the attention layer (0.18) indicates that the difference in behavior is more pronounced in the feedforward component.
These observations suggest a trade-off in how the network allocates its computational resources. The MAX+MED+SUM model appears to leverage its attention mechanism more, potentially to handle the diversity of operations it was trained on. In contrast, the SUM model seems to channel more of its computation through the feedforward network, possibly developing a more specialized but less flexible approach to solving the sum operation.
Figure 6: Attention patterns and layer dynamics in SUM vs MAX+MED+SUM (all3) models. Each panel shows (CoT) solution to a SUM modulo 10 problem, where ’>’ indicates solution steps. The first row shows the input sequence, with curved lines representing attention weights from Layer 2 in a 3-layer network. Black lines highlight attention patterns for a specific digit (shown in orange). Below are shown various layer metrics including the ratio of self-attention to feedforward norms (sa/ffwd), self-attention output norms (sa_out), feedforward output norms (ffwd_out), and ratios of layer outputs to inputs (sa_out/x, ffwd/x). Top: Model trained only on SUM operations shows attention primarily focused on parentheses and structural elements. Bottom: Model trained on MAX+MED+SUM (all3) shows attention strongly connecting to digits being combined in each CoT step, suggesting direct involvement in numerical computation. These distinct patterns suggest fundamentally different algorithms learned by each model.
Figure 7: Comparison of layer output/input ratio distributions for models trained on MAX+MED+SUM vs pure SUM. Left: Attention layer ratio histogram. Right: Feedforward layer ratio histogram. The x-axis represents the log ratio of layer output norm to input norm, while the y-axis shows the density. The black line represents the difference MAX+MED+SUM minus SUM distributions.
Supplemental Material
Appendix A Additional plots
A.1 Modulo 26
We define the following token vocabulary ABCDEFGHIJKLMNOPQRSTUVWXYZse()+-/%=>, where letters are mapped to integers such that A $\rightarrow$ 0, B $\rightarrow$ 1, …, Z $\rightarrow$ 25.
Figure 8: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 9: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 10: PCA of embeddings Odd vs Even: Same plot as above, only odd numbers colored red and even colored blue. Mixed SUM shows a clear odd-even separation in a few of the top PCs. Such a separation is not clearly observed in other cases. Interestingly, Pure SUM approximately separates odd-even, suggesting such separation may play a role in its algorithm.
A.2 Modulo 10
We conducted the same experiments also on mod 10. The smaller number of numbers makes definitive statements about some of the patterns more challenging. But all the patterns we observed in mod 26 also have parallels in mod 10, including the prominent odd-even split for operations involving SUM. Token vocabulary: %()+-/0123456789=>es
Figure 11: PCA of embeddings: We choose all models which reached over $90$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure MAX, or all mixtures involving a given operation, e.g. Mixed SUM. Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do. Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
Figure 12: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 13: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 14: Emergence of of abilities in ListOps, Mod 10: Each plot shows the same group of small transformer models trained on different variants of ListOps. Each variant uses a different mix of the four operations MAX, MIN, MED, SUM. Red dots are model reaching more than 50% accuracy, and blue are less than 50%. The dashed green line is a logistic fit and the yellow star indicates the transition point at 50%. The x axis is the model size and the plots are sorted in ascending order of transition points. The bottom is a bar plot showing the model size at the transition point. We observe that SUM is a clear outlier, with models requiring significantly more parameters to learn SUM. Surprisingly, combining SUM with other operations dramatically reduces the transition point, with model less than half the size easily reaching 100% accuracy.
Figure 15: Shuffled symmetric sum table, Mod 10: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version. Additionally, we observe that MAX+MED+Shuffled SUM is again more difficult than all operations except pure sum, suggesting that the number properties played an important role in the other tasks becoming easier in mixed training.
Figure 16: PCA of embeddings mod 10 for Shuffled symmetric SUM, pure vs mixed with MAX and MED: The numbers are colored based on parity (odd is red, even is blue). It is not expected that the embeddings show strong signals. There seems to be partial ordering of the numbers, but the clear wave patterns and clear parity separation is not evident.
Figure 17: PCA of embeddings with the same sample size (67): We choose the smallest 67 models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. We observe that the PCs get slightly distorted with the smaller sample size compared to fig. 13, but the overall structure stays the same. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 18: Parameters of continuous logistic regression fitting accuracy to number of parameters (No Early stopping).
Figure 19: Evolution of training loss, accuracy, and principal components of cosine similarities in the embedding layer. Modulo 10 ListOps. The top main figure shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 128 and 3 layers, trained either on SUM-only data (blue) or on mixed MAX+MED+SUM data (red). Curves represent the mean across three independent runs, with shaded regions indicating one standard deviation. All models were trained for 20000 iterations. The red and blue boxes beneath the main plot display the average embedding representations at different training stages (indicated by vertical gray dashed lines). PCA reveals that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss. In contrast, models trained solely on SUM data exhibit no clear structure in the embedding space and show long plateaus in the loss curve. This suggests that discovering an effective algorithm for the SUM operation in isolation requires significantly more exploration during training, in contrast to the more efficient joint training setting.
Figure 20: Evolution of training loss, accuracy, and PCs of cosine similarities in the embedding layer. ListOps base 10. The top main figure shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 128 and 3 layers, trained either on SUM-only data (blue) or on mixed MAX+MED+SUM data (red). Curves represent the mean across three independent runs, with shaded regions indicating one standard deviation. All models were trained for 20000 iterations. The red and blue boxes beneath the main plot display PCs revealing that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss. In contrast, models trained solely on SUM data exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
Appendix B Triplet Experiments
To increase control in our experimental setup, we train models only on triplet inputs 21. We construct 1000 unique triplet samples, splitting them into 900 for training and 100 for testing.
For joint training on the MAX, MED, and SUM operations, we use a combined training set of 2700 triplets (900 per operation) and evaluate performance on the same 100 excluded SUM samples. For a fair comparison, we also create a balanced training set of 900 samples by selecting 300 examples from each of MAX, MED, and SUM.
With early stopping, models trained on the full 2700-sample mixed dataset achieve 100% accuracy on the SUM test set, while those trained on only 900 SUM samples reach approximately 50% accuracy. When trained for 50k steps without early stopping, models eventually learn the task, but SUM requires significantly more parameters—consistent with earlier observations. These results highlight that learning the SUM operation in isolation is more challenging and slower (Fig. 21).
Figure 21: Training only on triplet ListOps data. The first row shows the same group of small transformer models trained on different variants of the triplet dataset. Each training set is constructed from 900 unique triplets. The 2700 MAX+MED+SUM dataset includes all 900 triplets, each labeled with three different operations. The 900 MAX+MED+SUM dataset contains 300 examples per operation, randomly sampled from the 2700 set. The 900 SUM dataset contains only the 900 unique SUM triplets. In all cases, the test set comprises the same held-out 100 SUM triplets. Red dots indicate models that exceed 50% test accuracy, while blue dots denote models that fall below this threshold. The dashed green line represents a logistic fit, and the yellow star marks the transition point at 50% accuracy. The x-axis shows the model size (in number of parameters), and subplots are ordered by increasing transition threshold. The red and blue boxes beneath the main plot display the average PCA embedding representations at selected regions; red corresponds to high-accuracy models, while blue indicates low-accuracy ones. Models trained on the combined MAX+MED+SUM task develop a discernible structure during training, unlike models trained solely on SUM. Notably, augmenting MAX and MED data with only 300 SUM samples allows the model to achieve 50% accuracy on a held-out 100 SUM sample test set, using a model 10× smaller than those trained on 900 SUM samples alone, which only reach $\approx 80\%$ accuracy.
Appendix C Related works
Scaling laws for language models characterize how validation loss scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design. However, this perspective overlooks critical aspects of learning. Our results show that the composition of the training data—specifically, the mix of tasks—can significantly alter what and how a model learns. Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
Joint training Joint training on compositional tasks allows models to acquire the foundational primitives necessary for solving complex operations, often leading to abrupt performance improvements once all constituent skills are learned lubana2024percolation; okawa2023compositional. Similarly, training on multiple arithmetic tasks has been shown to improve accuracy on individual operations, highlighting the benefits of shared representations for compositional generalization lee2024teaching.
Number representation Neural networks trained on modular addition tasks develop structured embedding spaces that reflect underlying arithmetic operations. Mechanistic analyses have shown that small models can exhibit ordered number patterns and implement distinct algorithmic strategies, depending on initialization and hyperparameters zhong2023clock. Some models converge to known solutions such as the Clock algorithm, while others discover novel procedures like the Pizza algorithm, illustrating the algorithmic diversity that can emerge from fixed training data. Periodic structures in the embeddings can be characterized via Fourier analysis, offering additional interpretability nanda2023progress. These behaviors have also been linked to grokking dynamics, where models abruptly generalize after extended training, accompanied by the emergence of structured embedding patterns power2022grokking; liu2022towards.
In general language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore it allows us to tune the task difficulty by combining different operations.
Here, we investigate how language models acquire arithmetic skills by training small models on the ListOps dataset, which enables explicit evaluation through structured mathematical expressions. Adopting a bottom-up approach, we find that models learn to solve these tasks once the number of trainable parameters surpasses a critical threshold. This threshold shifts with the difficulty of the target operation (MAX = MIN < MED < SUM), indicating a dependency on task difficulty. Surprisingly, joint training on multiple operations often facilitates learning, outperforming models trained on individual operations (e.g. MAX+MED+SUM < SUM). This suggests that task diversity can ease optimization by promoting shared representations.
Appendix D Data and processing
D.1 Dataset Notation
ListOps consists of nested mathematical equations involving operations such as min, max, median, and sum modulo 10 applied to single-digit numbers (0-9). It uses the Polish notation: (operation, [inputs]) For example: max(3,min(7,4,9))=4.
[…]
In this notation, ’s’ denotes the start of the expression, ’e’ marks the end, and parentheses indicate nesting levels.
D.2 Tokenization
We employ a character-based tokenization strategy for processing ListOps expressions. This approach offers several advantages:
  1. 1.
    Simplicity: Character-level tokenization eliminates the need for complex tokenization rules or a large vocabulary.
  2. 2.
    Generalizability: It allows the model to potentially generalize to unseen number combinations or deeper nesting levels.
Each character in the ListOps expression, including digits, operation symbols, and structural elements (parentheses, ’s’, ’e’), is treated as a separate token. This granular representation enables the model to learn the syntactic structure of the expressions alongside their semantic content.
D.3 Chain of Thought Implementation
We find that directly solving nested ListOps in one step can be quite challenging for transformer model (Fig. 22) Even with a maximum of three nesting levels with three operands (inputs) we find that GPT models with over 10 million parameters still fail to learn the task. To enhance model performance, particularly on more complex operations like sum modulo 10, we introduced a chain of thought (CoT) approach in our training data. This method involves providing step-by-step solutions that resolve the deepest nesting level at each step. For example:
s(%12(%34))>(%127)>0=0e
In this CoT representation:
  • The initial expression is s(%12(%34))
  • The first step resolves the innermost operation: (%34) becomes ‘7‘
  • The intermediate result is shown: s(%12(7))>(%127)
  • The process continues until the final result is reached: s(%12(7))>(%127)>0=0e
This CoT approach serves multiple purposes: 1. It guides the model through the problem-solving process, mimicking human-like reasoning. 2. It provides more granular supervision, potentially aiding in learning complex operations. 3. It allows us to study how models learn to break down and solve nested problems. Our experiments show that this CoT method significantly improves model performance, particularly for the challenging sum modulo 10 operation (Fig. 22).
Figure 22: no - CoT vs. CoT. Providing solutions as chain-of-thought (CoT) helps models learn the tasks. In almost all cases, CoT accelerates learning, enabling smaller models to succeed. This effect is especially strong for the sum operation, which cannot be learned without CoT.
Appendix E Observations from the norm of Attention and Feedforward outputs
To investigate the internal dynamics of our models, we focused on the final layer of a 3-layer transformer network, featuring a single attention head and an embedding dimension of 128. Our analysis centered on comparing the behavior of models trained on ”ALL3” operations (MAX+MED+SUM) versus those trained solely on the ”SUM” operation.
We introduced a novel metric to quantify the impact of different components within the network: the ratio of output to input norms for both the self-attention (SA) and feedforward (FFN) sublayers. Specifically, we computed:
  1. 1.
    Attention ratio: $r_{attn}=\frac{\|SA(LN_{1}(x))\|}{\|x\|}$
  2. 2.
    Feedforward ratio: $r_{ffwd}=\frac{\|FFN(LN_{2}(x_{1}))\|}{\|x_{1}\|}$
where $LN_{1}$ and $LN_{2}$ are layer normalization operations, and $x_{1}$ is the output of the self-attention sublayer. These ratios provide insight into how much each component modifies its input, serving as a proxy for the component’s impact on the overall computation.
We analyzed the distribution of these ratios across a test set consisting of sum operations for both the ’ALL3’ and ’SUM’ models. Kernel Density Estimation (KDE) plots were used to visualize the distributions, and we employed several statistical measures to quantify the differences.
Attention Sublayer
The attention sublayer showed moderate but statistically significant differences between the ’ALL3’ and ’SUM’ models:
  • Kolmogorov-Smirnov test: statistic = 0.1592, p-value < 0.0001
  • Jensen-Shannon divergence: 0.1591
  • Wasserstein distance: 0.0696
  • Effect size (Cohen’s d): 0.1761
  • 95% CI for mean difference: (0.0466, 0.0860)
The KDE plot revealed that the ’ALL3’ model’s attention ratio distribution was more concentrated and peaked higher than the ’SUM’ model’s distribution. The positive effect size and confidence interval indicate that the ’ALL3’ model generally had higher attention ratios.
Feedforward Sublayer
The feedforward sublayer exhibited more pronounced differences:
  • Kolmogorov-Smirnov test: statistic = 0.2461, p-value < 0.0001
  • Jensen-Shannon divergence: 0.1617
  • Wasserstein distance: 0.2830
  • Effect size (Cohen’s d): -0.3379
  • 95% CI for mean difference: (-0.3042, -0.2226)
The KDE plot for the feedforward ratios showed a clear shift between the two distributions. The ’SUM’ model’s distribution was shifted towards higher values, as confirmed by the negative effect size and confidence interval.
Interpretation
These results reveal distinct operational patterns between models trained on ’ALL3’ operations versus those trained solely on ’SUM’:
  1. 1.
    In the attention sublayer, the ’ALL3’ model shows slightly higher ratio values, suggesting that attention mechanisms play a more pronounced role when the model is trained on diverse operations. This could indicate that the attention sublayer is capturing more complex patterns or relationships necessary for handling multiple operations.
  2. 2.
    Conversely, in the feedforward sublayer, the ’SUM’ model demonstrates significantly higher ratio values. This suggests that when trained on ’Sum’ alone, the model relies more heavily on the feedforward network for computation. This could imply that the ’SUM’ operation is being implemented more directly through feedforward transformations.
  3. 3.
    The larger effect size in the feedforward layer (-0.3379) compared to the attention layer (0.1761) indicates that the difference in behavior is more pronounced in the feedforward component.
These observations suggest a trade-off in how the network allocates its computational resources. The ’ALL3’ model appears to leverage its attention mechanism more, potentially to handle the diversity of operations it was trained on. In contrast, the ’SUM’ model seems to channel more of its computation through the feedforward network, possibly developing a more specialized but less flexible approach to solving the sum operation.
This analysis provides evidence that the internal dynamics of transformer models adapt significantly based on the diversity of tasks they are trained on, even when evaluated on the same type of operation (’SUM’). It highlights the importance of considering task diversity in understanding and optimizing neural network architectures.
Figure 23: Comparison of layer output/input ratio distributions for models trained on all three operations (all3 - MAX+MED+SUM) versus sum operation alone (SUM). Left: Attention layer ratio histogram. Right: Feedforward layer ratio histogram. The x-axis represents the log ratio of layer output norm to input norm, while the y-axis shows the density. The black line represents the difference between the all3 and sum distributions (all3 - sum). These plots illustrate distinct operational patterns between the two models, with the attention layer showing increased activity in the all3 model and the feedforward layer demonstrating higher ratios in the sum model.
Appendix F Ablation studies
Figure 24: Random sum table. Joint training helps in learning even randomized sum tables. The numbers in the sum table follow a uniform distribution. We found that a randomized sum table with a non-uniform distribution is also easy to learn. However, a randomized sum table with the same uniform distribution remains as hard to learn as the original sum table. Early stopping criteria were applied during training in all simulations.
Figure 25: Whittaker-Shannon number representation for cyclic SUM Modulo 10. Top row: Cyclic matrices generated using Whittaker–Shannon interpolation illustrate how digit embeddings can encode modular addition via matrix multiplication. Middle row (red box): Embeddings learned by a model trained on MAX+MED+SUM. Bottom row (blue box): Cyclic SUM representation constructed using Whittaker–Shannon numbers and digit ordering, showing a similar structure to the model’s embeddings—suggesting that the model discovers an emergent encoding of cyclic SUM.
Appendix G Complexity: Number of operands and nesting level.
In previous studies, task complexity has been characterized through various measures, including the number of bits required to memorize the task, which corresponds to the length of the expression dave2024investigating, the number of operands and nesting depth petruzzellis2024benchmarking, and the structure of the computational graph dziri2024faith, which captures the number of nestings. Here, we examine how the number of operands and nesting levels affect the learning ability of small language models. Focusing on the all3 task, which combines max, med, and sum operations, we manipulate complexity by varying the number of operands (arg = {3, 4, 5}) and nesting depth (depth = {3, 4, 5}). The length of equations, measured by the number of characters, depends on both the number of operands and the nesting level. We find that nesting has a greater impact: increasing the nesting level from 3 to 4 results in longer equations than increasing the number of operands from 3 to 5 at a fixed nesting level. We train the GPT model on the all3 dataset with all combinations of arg (number of operands) and depth (nesting levels), finding that the model’s performance correlates with the sum of operands and nesting levels (arg + depth). Notably, transition points tend to group together for configurations with the same sum (Fig. 26, 27b). While Fig. 27a demonstrates that the model requires more parameters to solve longer equations, it also indicates that arg + depth serves as a reliable predictor of the transition point.
Figure 26: Learning MAX+MED+SUM operations with varying numbers of operands and nesting levels. The model requires more parameters as the number of operands and nesting levels increases. Higher nesting levels particularly demand larger model sizes to learn the task. We present the average of five simulations for each configuration and fit a sigmoid function, with the cross marking the middle value (transition point). The transition points reveal an interesting pattern: the sum of the number of operands and nesting levels groups together. For example, the transition points for arg-3,depth-4 (orange) and arg-4,depth-3 (red) are close to each other, as are those for arg-4,depth-5 (brown), and arg-5,depth-4 (grey). Early stopping criteria were applied during training in all simulations.
(a) Average equation length vs. transition point.
(b) Transition points.
Figure 27: Transition point vs. Equation Lenght and Nesting Depth. (a) Transition point in function of the average equation length. (b) Heat plot of transition point in function number of operands and nesting depth.
Appendix H Experiment settings.
  1. 1.
    Model: GPT model (karpathy2022nanogpt)
  2. 2.
    Number of layers: 1, 2, 3, 4, 5, 6
  3. 3.
    Number of head: 1
  4. 4.
    Embedding dimension: 4, 8, 12, 16, 24, 32, 48, 64, 96, 128, 192, 256
  5. 5.
    Number of random seed: 5
  6. 6.
    Context window size: 128
  7. 7.
    Batch size: 64
  8. 8.
    Optimizer: Adam
  9. 9.
    Minimum learning rate: 1e-4
  10. 10.
    Maximum interation steps: 20000/50000
  11. 11.
    Early Stopping criteria: Early stopping was applied after 2000 iterations when the change in training loss was below $\Delta_{min}$ = 2.5e-4 for 10 consecutive evaluation steps.
  12. 12.
    Data We generate 50,000 initial equations and split them into training and test sets by excluding 100 randomly selected triplets. The training set consists of approximately 45,000 examples that do not contain any of the excluded triplets. The test set comprises around 2,000 examples, carefully curated to ensure that the excluded triplets do not appear, even in the final step of each equation. We evaluate model performance on a subset of 1,000 final test examples.
  13. 13.
    Vocab - Base 10 ListOps: %()+-/0123456789=>es
  14. 14.
    Vocab - Base 26 ListOps: ABCDEFGHIJKLMNOPQRSTUVWXYZse()+-/%=>
  15. 15.
    Hardware: All simulations were run on a mix of GPUs, including NVIDIA A100, H200, RTX 4090, and A30, as well as on a standard modern laptop. The experiments are lightweight, requiring approximately 1.5GB of RAM, and can be executed efficiently on any recent laptop-class device.
Appendix I Experiments on ListOps.
  1. 1.
    Base 10: - no CoT/CoT: MAX, MED, SUM, MAX/MED/SUM
  2. 2.
    Base 10: - all combination: MAX, MIN, MED, SUM + MAX/MED/SUM-RANDOM SHUFFLED
  3. 3.
    Base 26: - all combination: MAX, MIN, MED, SUM + MAX/MED/SUM-RANDOM SHUFFLED
  4. 4.
    Triplet: MAX+MED+SUM 2700 (2700 training data samples = 900 MAX + 900 MED + 900 SUM), MAX+MED+SUM 900 (900 training data samples = 300 MAX + 300 MED + 300 SUM), SUM (900 training data samples)
Appendix J Shuffled SUM modulo 26
We make a symmetric sum table ($A+B=B+A$), with a randomly shuffled right hand side, meaning where if in the table $A+B=C$, $C$ does not the actual arithmetic sum of $A+B$ modulo 26. We do this to test a couple of hypotheses:
  1. 1.
    Hypothesis 1: The SUM trained alone is not really learning the logic arithmetic of numbers, but rather memorizing the sum table.
  2. 2.
    Hypothesis 2: Joint training with max and med leads to learning number properties.
If the first hypothesis holds, the shuffled sum would also require similarly high number of parameters as the normal sum and show similarly random patterns in the embedding space. If the second hypothesis holds, then joint training of shuffled sum should actually have detrimental effects on learning sum because number properties don’t play a role in learning shuffled sum. Thus we expect the jointly trained model to struggle to learn all three operations, or require significantly higher number of parameters to master shuffled sum, compared to joint training on regular sum.
Figure 28: Shuffled symmetric sum table, Mod 26, 20k steps: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version. We also observe that MAX+MED+Shuffled SUM never reaches more than 80% accuracy. The third scatter plot from the let shows the accuracy of the MAX+MED+Shuffled SUM model on the Shuffled SUM test set. We see that the accuracy is very low ($\sim 20\%$ top), showing that the mixed model never learned the shuffled SUM. This may suggest that MAX+MED revealed number properties, but Shuffled SUM was incompatible with those properties, leading to a model that overall cannot solve the two problems (MAX+MED and Shuffled SUM) simultaneously.
Figure 29: Shuffled symmetric sum table, Mod 26, 40 steps: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version. We also observe that MAX+MED+Shuffled SUM never reaches more than 80% accuracy. The third scatter plot from the let shows the accuracy of the MAX+MED+Shuffled SUM model on the Shuffled SUM test set. We see that the accuracy is very low ($\sim 20\%$ top), showing that the mixed model never learned the shuffled SUM. This may suggest that MAX+MED revealed number properties, but Shuffled SUM was incompatible with those properties, leading to a model that overall cannot solve the two problems (MAX+MED and Shuffled SUM) simultaneously.
Figure 30: PCA of embeddings for Shuffled symmetric SUM, pure vs mixed with MAX and MED, Mod 26: The numbers are colored based on parity (odd is red, even is blue). There seems to be partial ordering of the numbers, but the clear wave patterns are not visible. We do almost observe parity separation in PC3 and PC4, albeit with some noise. It is curious that the system still learns partial parity, but evidently this feature did not allow the system to learn the shuffled SUM perfectly.
Figure 31: Emergence of abilities in ListOps, 40k steps: Long training. Each plot shows the same group of small transformer models trained on a different mix of the four operations MAX, MIN, MED, and SUM. Red dots are models reaching more than 50% accuracy, and blue dots are less than 50%. The dashed green line is a logistic fit, and the yellow star indicates the transition point at 50%. The x-axis is the model size (number of parameters), and the plots are sorted in ascending order of transition points. The bottom panel shows a bar plot of the model sizes at the transition points, with each group distinguished by a different color.
Appendix K Attention and Feedforward outputs modulo 10
Figure 32: Attention layer and Feedforward layer ratio histogram. Each row shows the attention and feedforward layer ratio histogram for models trained on MAX, MED, SUM, MAX+MED,and MAX+MED+SUM. Each column shows the ratio histogram in different attention blocks. The model had 128 embedding and 3 layers and all models rechec 99% accuracy.
Figure 33: The mean attention layer ratio vs the feedforward layer ratio. Each plots show the means in different layers.
Figure 34: Comparison of layer output/input ratio distributions for models trained on all three operations (MAX+MED+SUM) versus sum operation alone (SUM). The title of the figures contain the Effective size showing the difference between the model.

### v2

@@
-The Role of Task Complexity in Emergent Abilities of Small Language Models
+Small Models, Smarter Learning: The Power of Joint Task Training
 Abstract
-We investigate the relationship between task complexity and the minimum model size required for learning specific tasks in small transformer models.
-We focus on the ListOps dataset, consisting of nested math operations.
-We define the task complexity as the Kolmogorov complexity (KC) of the code solving the task, using a rough proxy for KC.
-We find a power-law relation between KC and parameters required to learn, suggesting number of parameters to learn harder task increases almost cubic in KC.
-On individual math operations, sum mod 10 is hardest to learn.
-Surprisingly, when combining tasks, we observe that sum is learned earlier and with fewer parameters when trained alongside max and median.
-Analyzing the model, we find strong evidence that models trained on sum alone and models trained jointly converge to different algorithms.
-Concretely, the sum alone model doesn’t seem to have learned number properties in the embedding layer, likely memorizing the sum table.
-In contrast, models trained on three tasks (maximum, median and sum) reveals that joint training results in clear number-like properties.
-Finally, we also find evidence that the sum-only model utilizes its feedforward layer more than the jointly trained model.
-Conversely, the attention layer in the joint model is activated more than the sum model.
-Our findings suggest there is another dimension to emergent abilities in language models, namely the algorithms being learned, potentially impacting scaling laws.
 […]
-1 Related works
-Scaling laws for language models characterize how validation loss scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design.
+Machine Learning, ICML
+1 Introduction
+Scaling laws predict language model performance as a function of model size (Hoffmann et al., 2022), compute (Muennighoff et al., 2023), and dataset size (Hestness et al., 2017), but treat training data as homogeneous and do not explain how task composition shapes learning efficiency.
+Prior work has shown that multi-task learning improves generalization (Caruana, 1997; Ruder, 2017) and that joint arithmetic training can benefit individual tasks (Lee et al., 2024; Abedsoltan et al., 2025).
+However, a systematic understanding of when joint training helps versus fails—and by how much—remains lacking.
+[…]
+[…]
+Mathematical tasks provide an ideal testbed for this investigation: they allow precise control over task difficulty, quantitative evaluation of accuracy, and systematic combination of operations (Nangia and Bowman, 2018; Nye et al., 2021; Dziri et al., 2023).
+We use the ListOps dataset (Nangia and Bowman, 2018), which consists of nested operations such as MAX(2,3,ADD(4,5)), extended with modular arithmetic and permutation groups.
+Summary of contributions.
+We study how small transformer models learn nested operations—maximum (MAX), minimum (MIN), median (MED), modular addition (ADD), modular product (PROD), and alternating addition-subtraction (NADD: $\sum_{i}(-1)^{i}x_{i}$)—both in isolation and jointly.
+We further extend to permutation groups, where products can be decomposed into sub-block operations.
+Our main findings are:
+  1. 1.
+    Task difficulty hierarchy: MAX/MIN are easiest; MED and PROD (non-prime modulus) are intermediate; ADD, NADD, and PROD (prime modulus) are hardest, requiring an order of magnitude more parameters.
+  2. 2.
+[…]
+  3. 3.
+    Synergies that fail: Pairing structurally similar hard tasks (ADD+NADD) provides no benefit.
+Nor does pairing ADD with PROD when the modulus is prime.
+This indicates that task synergy depends on shared computational primitives, not superficial similarity.
+  4. 4.
+    Mechanistic evidence: PCA of learned embeddings shows that successful joint training induces structured number representations (ordering, parity, divisibility) absent in single-task models.
+  5. 5.
+[…]
+These results demonstrate that joint training can guide models toward more efficient solutions, revealing structure in scaling laws that depends critically on task composition.
+Understanding these synergies has practical implications: if certain task combinations make models “smarter” by inducing better representations, this could inform curriculum design for pretraining—a form of pre-pretraining that reduces the resources needed for downstream tasks.
+2 Related Work
+[…]
+Curriculum learning.
+Presenting tasks in order of difficulty improves learning (Bengio et al., 2009; Yin et al., 2024; Kim and Lee, 2024).
+Our transfer experiments (Section 4.3) relate to curriculum learning, showing that pretraining on easy tasks enables learning hard tasks at smaller model sizes.
+However, our focus is on simultaneous joint training rather than sequential curricula.
+Grokking and mechanistic interpretability.
+Work on grokking (Power et al., 2022; Liu et al., 2022) and mechanistic interpretability (Nanda et al., 2023) has shown that models solving modular arithmetic develop structured internal representations.
+We build on these insights by using embedding analysis to understand why joint training helps: successful task pairings induce structured embeddings (ordering, parity) that single-task training does not.
+Our setup differs from grokking studies in that we vary model size at fixed dataset size, rather than observing delayed generalization at fixed model size.
+We provide the first systematic map of task compatibility for arithmetic operations, extending to permutation groups and confirming robustness to architectural choices (Appendix H).
+3 Methodology
+We train small-scale transformer models and analyze both their performance and internal representations.
+This section outlines our methodology, detailing the dataset, model architecture, and evaluation protocols.
+Choice of task.
+ListOps consists of nested mathematical equations involving operations.
+Our setup uses MAX, MIN, MED (median), ADD (addition modulo $n$), PROD (product modulo $n$) and NADD (alternating addition and subtraction modulo $n$) applied to numbers (0 to n1n-1).
+We conduct experiments using $n=10$ to $n=226$ (Appendix F.3.1).
+ListOps provides procedural generation with controlled difficulty, exact evaluation, adjustable complexity via nesting depth, and systematic investigation of inter-task relationships.
+Dataset description.
+We use a simplified functional notation of the form $f(x,y,\ldots)$.
+For example, max_10(3,min_10(7,4,9))=4<eos>, where <eos> denotes the end token.
+The subscript in the function name indicates the modulo operation e.g., max_10 represents the max function modulo $10$.
+Tokenization and CoT.
+We use word-level tokenization: each operation (e.g., add_20), each number (e.g., 13), and each symbol ((, ), =, <eos>) is a single token (Appendix F.3.2).
+Since directly solving nested expressions proves challenging (Fig.
+24), we use chain-of-thought (CoT) that decomposes expressions into intermediate steps (Appendix F.3.3): add_10(9 add_10(3 4))>add_10(9 7)>6=6<eos>, where > indicates solving the innermost operation.
+This follows the scratchpad approach of Lee et al.
+(2024).
+Task difficulty increases with the number of operands and nesting depth (Appendix Fig.
+48).
+Train-Test split.
+The experiments were conducted on data generated with a maximum nesting depth of 2 and up to 3 operands per operation.
+Equations are constructed from digit sequences forming duplets and triplets (e.g., ”1,0”, ”0,0,1”).
+Since the CoT format can produce many intermediate patterns within a single equation, we implement a rigorous train-test split to prevent data leakage: we designate 100 of the 1000 possible triplets (e.g.
+”749”) as an exclusion set.
+Training equations never contain any excluded triplet, while test equations each contain at least one excluded triplet, ensuring that test set patterns are novel.
+Permutation groups.
+To test generality beyond arithmetic, we extend ListOps to permutation groups.
+By Cayley’s theorem, every finite group is isomorphic to a permutation subgroup, making this a natural generalization.
+We construct $6\times 6$ block-diagonal permutation matrices (two $3\times 3$ blocks) and define three operations: OP (full matrix, 36 elements), OP_TOP and OP_BOTTOM (individual blocks, 6 elements each).
+See Fig.
+9a and Appendix Figs.
+1513 for details.
+Performance evaluation.
+We randomly select 1000 equations from the held-out test set.
+The model is prompted to generate solutions character-by-character, starting from the equation prompt (e.g., add_10(3,min_10(7,4,9))>).
+We use the output to produce two metrics: Loss: using cross-entropy computed for every character of the output.
+Accuracy: based on the number of correct answers evaluated using only the final answer, which we define as the first character after the first ‘==’ symbol.
+Figure 2: Joint training with compatible tasks reduces the model size needed to learn.
+Each panel shows accuracy vs.
+model size for operations trained in isolation ($\bullet$), with an easier task ($\blacksquare$), or with a harder task ($\blacktriangle$).
+Left: ADD benefits from PROD ($2.5\times$ smaller models) but not from NADD.
+Center: NADD benefits from PROD but not from ADD.
+Right: PROD benefits from MAX ($2.6\times$ smaller) but not from ADD+MAX.
+Dashed lines show logistic fits; legends report the 80% accuracy transition point.
+Model architecture.
+For our experiments, we employed a series of tiny GPT models, inspired by the nanoGPT architecture (Karpathy, 2022).
+Unlike the standard sequential transformer, we implemented a recurrent variant in which a single transformer block was iteratively applied by feeding back its output as the next input (see the architecture Appendix Fig.
+23).
+This recurrent design improved learning efficiency and yielded more structured embedding patterns (see results from traditional deep transformer models in the Appendix F).
+Each model in our study uses a single attention head.
+We set the feedforward hidden dimension to four times the embedding dimension, a common practice in transformer architectures, providing sufficient complexity in the feedforward networks while keeping the model size constrained.
+By varying the embedding size we create a range of models with different parameter counts.
+Ablation studies with 1–4 attention heads show no qualitative change in results (Appendix H), and experiments with standard deep transformers (2–4 layers) yield similar patterns, though with less structured embeddings (Appendix F).
+Learning transition.
+We define the learning transition as the minimum model size at which a task can be learned.
+Operationally, we fit a logistic function to accuracy vs.
+model size and define the transition point as the model size reaching 80% of the asymptotic accuracy.
+We chose 80% rather than 50% because the latter corresponds to models that have not yet reliably mastered the task; 80% ensures the model has learned a generalizable solution rather than partially memorizing patterns.
+The transition curves are remarkably consistent across runs for the recurrent architecture with fixed dataset size.
+Dataset scaling.
+[…]
+4 Results
+Figure 3: Shuffled ADD: joint training fails when number structure is destroyed.
+Left: Learning curves for various tasks including Shuffled ADD (randomly permuted addition table).
+Right: Transition points—Shuffled ADD is nearly as hard as pure ADD.
+Unlike regular ADD, joint training with MAX+MED hurts Shuffled ADD: the mixed model reaches only $\sim$20% accuracy (third panel).
+This confirms that MAX+MED induces number-like embeddings incompatible with the randomized table.
+To understand how language models acquire mathematical abilities, we focus on accuracy as the primary metric for observing the emergence of learning, following the approach of the emergent abilities literature (Wei et al., 2022).
+While some subsequent studies have questioned the concept of emergence by examining other metrics (Schaeffer et al., 2023), we argue that this critique overlooks similar patterns observed in physics: in a phase transition, only some quantities may exhibit discontinuous changes.
+A metric that can capture the emergence of the new phase is called the “order parameter”.
+Hence, we choose to use accuracy as our order parameter.
+Experiments use a recurrent transformer with embedding dimension $n_{embed}$ ranging from 8 to 362, and 4 recurrence steps (we have found that using 4 recurrence steps yields the best performance).
+Similar to the emergence literature, we observe the total number of parameters to be the strongest indicator of accuracy, rather than $n_{embed}$ or depth.
+Our experimental design focuses on identifying the minimum model size sufficient for learning each task or task combination—once this threshold is exceeded, larger models reliably reach high accuracy, making further scaling unnecessary for our analysis.
+We primarily present results for modulo 20, with similar patterns observed for other modulos (Appendix F.5 - mod $10$ results, F.4 - mod $26$ results).
+Experiments with higher moduli reveal a basic scaling relationship: larger moduli require proportionally larger models and more training data to achieve comparable performance.
+We find that joint training on compositional tasks substantially improves performance and reduces the parameter requirements for learning individual operations.
+Training on multiple arithmetic tasks helps models learn the basic building blocks needed for solving more complex problems.
+This shared representation not only accelerates learning but also produces abrupt transitions in performance once the necessary constituent skills are mastered (Lubana et al., 2024; Okawa et al., 2023).
+However, joint training is not universally beneficial, and task compatibility is more nuanced than surface-level similarity suggests.
+Our experiments reveal that effective transfer requires overlap in underlying computational primitives and mathematical structures, not merely similar algorithms or comparable difficulty.
+Pairing hard problems together often increases required model size when tasks lack shared low-level building blocks, demonstrating that joint training can actively interfere with learning rather than facilitate it.
+These findings extend previous work on compositional generalization (Lee et al., 2024) by systematically characterizing when multi-task learning helps or hinders algorithmic task performance, showing that compatibility depends on deep structural properties rather than intuitive task similarity.
+Transition to learning.
+Figure 2 shows learning transitions for single and joint operations.
+We find that joint training enables models to solve more difficult tasks with fewer parameters and less data.
+For example, ADD and NADD are the most demanding tasks when trained in isolation, yet when combined with PROD, models that are $\times$2.5 smaller are able to reach 100% accuracy on the task.
+Similarly, while PROD on prime moduli is challenging on its own, pairing it with MAX reduces the required model size by a factor of $\times$2.6.
+In contrast, pairing ADD with NADD does not decrease the required model size, nor does pairing PROD-prime with ADD or NADD.
+These results challenge the assumption that harder tasks always demand larger models.
+Joint training helps only when tasks share the right structure: ADD and NADD use similar algorithms but require different representations, so pairing them fails.
+Meanwhile, PROD with non-prime modulus helps ADD because divisibility patterns create shared computational primitives.
+The key insight: task synergy requires overlapping low-level building blocks, not superficial similarity.
+These patterns extend to combinations of 3–4 tasks: joint training on all combinations of MAX, MIN, MED, and ADD (modulo 26) shows the same grouping of transition points by task compatibility (Appendix Fig.
+29).
+Testing the number properties hypothesis.
+We hypothesize that joint training helps because easier tasks reveal number properties (e.g., ordering, parity) that benefit harder tasks.
+To test this, we designed an experiment that destroys number relationships: we shuffle the addition table so that $A+B=C$ becomes a random mapping with no arithmetic structure (Appendix D).
+If joint training benefits ADD by revealing number properties, then Shuffled ADD should not benefit from joint training with MAX and MED.
+We conducted experiments using a shuffled, symmetric ADD table ($A+B=B+A$) in modulo 26.
+The shuffled addition proved more difficult to learn than all other mathematical operations except the original ADD (Fig.
+3).
+Crucially, MAX+MED+Shuffled ADD was harder than Shuffled ADD alone: the mixed model reached only 60% accuracy on Shuffled ADD and required twice as many parameters.
+This is the opposite of what we observe with regular ADD, where joint training helps dramatically.
+This confirms our hypothesis: MAX and MED induce number-like embeddings with ordering structure, but Shuffled ADD is incompatible with these representations.
+PCA analysis shows Shuffled ADD embeddings lack parity separation or ordering (Appendix Fig.
+22).
+4.1 Embedding Analysis: What Number Properties Emerge?
+The Shuffled ADD experiment confirms that number properties are key to joint training benefits.
+We now examine what specific properties emerge in the learned embeddings.
+To do this, we analyze the embedding layer using principal component analysis (PCA).
+PCA methodology.
+In our tokenization scheme, each number and operation is a single token.
+While embedding layers of different-sized models cannot be directly compared (they have different dimensions), we can compare their $n_{\text{vocab}}\times n_{\text{vocab}}$ cosine similarity matrices.
+To reduce noise, we average similarity matrices across models that learned each task (accuracy $90\%$), then compute PCA on the averaged matrix.
+This reveals consistent patterns across model sizes (see Appendices F.4 and F.5 for additional results).
+Structured embeddings in joint training.
+Figure 4 shows that joint training induces structured embeddings: PROD shows strong parity separation and ordering; ADD+PROD preserves this structure; pure ADD shows weak structure; NADD develops different structure; and ADD+NADD loses structure from both tasks.
+This explains why ADD+NADD fails despite algorithmic similarity: incompatible representations.
+Figure 4: PCA of embeddings reveals structured number representations in joint training.
+Top 5 principal components of the cosine similarity matrix, colored by parity (yellow=odd, purple=even).
+PROD: Strong parity separation and ordering structure.
+ADD+PROD: Preserves PROD’s structure, explaining why joint training helps ADD.
+Pure ADD: Weak, noisy structure.
+NADD: Highly structured but different from ADD.
+ADD+NADD: Loses structure from both tasks, explaining why this pairing fails.
+Quantifying embedding structure: separation score.
+Beyond parity, we observe that embeddings cluster numbers by their remainders modulo divisors of $n$.
+To quantify this, we define a separation score: for each candidate divisor $d$, we group numbers by their remainder mod $d$, compute centroids for each group in the top PCs, and measure the ratio of between-group distance to within-group standard deviation.
+High separation scores indicate strong clustering by that divisor (Fig.
+5).
+The results reveal: PROD embeddings cluster by divisors of $n$ (e.g., mod-2, 3, 6 for nn=42); prime moduli show no clustering, explaining why PROD-prime is hard; and ADD+PROD inherits PROD’s divisor structure while pure ADD shows weak clustering.
+This explains why PROD helps ADD only for non-prime moduli: divisibility structure creates transferable primitives.
+Figure 5: Embeddings cluster by divisors of the modulus.
+(a) PCA of PROD-45+ADD-45 embeddings colored by remainder class shows clear mod-3 and mod-5 clustering (divisors of 45).
+(b) Separation score (between-centroid distance / within-group variance) quantifies clustering strength across moduli.
+Non-prime moduli (42, 45) show strong clustering by their divisors; prime modulus (43) shows none.
+ADD+PROD inherits PROD’s divisibility structure; pure ADD shows weak clustering.
+This explains why PROD helps ADD only for non-prime moduli.
+4.2 Learning Dynamics: How Structure Emerges
+The embedding analysis shows that joint training produces structured representations.
+But when does this structure emerge during training?
+We examine the learning dynamics to understand how easy tasks scaffold hard tasks.
+Figure 6: Joint training improves training speed by encouraging structured embeddings during learning dynamics.
+(Top) Two models are trained: one on a joint MAX+MED+ADD dataset and one on a pure ADD dataset (mod 26).
+We observe that both training loss (solid) and accuracies (dashed) improve faster for the joint training model across all evaluated tasks (mean $\pm\sigma$ over 3 runs.).
+In joint training, MAX learns first, then MED, then ADD learns only after the embeddings become structured.
+(Bottom) PCA snapshots are taken during training (number parity is colored red/blue).
+Joint training accuracies rise quickly after the embeddings learn structure, which only happens in the joint training model.
+The pure ADD model shows no embedding structure and exhibits loss plateaus, suggesting it struggles to discover an efficient algorithm.
+Figure 6 compares pure ADD vs.
+MAX+MED+ADD ($n_{\text{embed}}=96$).
+In joint training, MAX learns first, then MED, then ADD—and structured embeddings emerge before ADD accuracy rises, suggesting structure enables learning.
+Joint training converges in half the steps; pure ADD shows no structure and loss plateaus.
+4.3 Causal Test: Can Learned Structure Transfer?
+To establish causation, we test whether structure from easy tasks transfers to hard tasks.
+We design a curriculum (Fig.
+7): train on ADD+PROD, gradually shift to pure ADD, and compare to a baseline trained only on ADD.
+The results are striking: baseline ADD stays at $\sim$15% accuracy, but curriculum achieves >>95%. Structured embeddings develop during ADD+PROD and persist through the transition.
+Figure 7: Curriculum learning enables transfer of structured representations.
+Left: PCA of embeddings at 0K, 10K, 20K, and 30K steps (top), colored by parity (red=odd, blue=even).
+Structure emerges during joint ADD+PROD training (mix ratio rises at 10K) and persists as the curriculum shifts to pure ADD.
+Loss and accuracy curves (bottom) show the curriculum transition.
+Right: Baseline ADD stays at $\sim$15% accuracy; curriculum reaches >>95%.
+Shaded: $\pm 1\sigma$ over runs.
+[…]
+4.4 Scaling Across Moduli
+Do the benefits (and failures) of joint training persist across different moduli?
+We systematically measure transition points for various task combinations across moduli ranging from 10 to 80 (Fig.
+8).
+The results cleanly separate into cases where joint training helps versus cases where it does not.
+Figure 8: Joint training helps only with compatible tasks when scaling across moduli.
+Top (helps): ADD+PROD (left) reduces transition points by $3.6\pm 1.6\times$ for non-prime moduli, while MAX+PROD (right) reduces PROD transition by $3.3\pm 1.3\times$ even for prime moduli.
+Bottom (no benefit): ADD+PROD (left) provides no benefit for prime moduli (ratio $0.85\pm 0.25\times$), while ADD+MAX+PROD (right) increases PROD-prime transition (ratio $0.77\pm 0.13\times$).
+Modulus values are indicated on the x-axes and as numbers on each scattered point.
+Joint training helps:
+ADD+PROD reduces transitions by $3.6\pm 1.6\times$ for non-prime moduli; MAX+PROD reduces PROD transitions by $3.3\pm 1.3\times$ even for primes (MAX provides ordering structure).
+Joint training fails:
+ADD+PROD provides no benefit for prime moduli ($0.85\pm 0.25\times$); ADD+MAX+PROD actually hurts PROD-prime ($0.77\pm 0.13\times$). These patterns hold robustly across moduli (Appendix Fig. 44).
+4.5 Generalization Beyond Arithmetic: Permutation Groups
+To test generality, we extend ListOps to permutation groups with no numerical structure.
+We construct $6\times 6$ block-diagonal permutation matrices (two $3\times 3$ blocks) with three operations: OP (full matrix, 36 elements), OP_TOP and OP_BOTTOM (individual blocks, 6 elements).
+The sub-block operations decompose the full operation into simpler components—analogous to how MAX helps ADD.
+Figure 9: Joint training benefits generalize to permutation groups.
+(a) Setup: $6\times 6$ block-diagonal permutation matrices with two $3\times 3$ blocks.
+OP acts on the full matrix (36 elements); OP_TOP and OP_BOTTOM act on individual blocks (6 elements each).
+(b) Transition curves: joint training vs.
+pure OP.
+(c) Sub-block operations reduce the OP transition point by $5\text{--}6\times$, confirming that decomposing complex operations into simpler primitives helps learning—even in non-numerical domains.
+[…]
+5 Discussion
+Why does joint training help?
+[…]
+Toward a predictive theory: output distributions.
+We hypothesize that task difficulty relates to the output distribution of $\text{Op}(a,b)=c$: tasks with uniform output distributions (like ADD) offer no patterns to exploit, while tasks with peaked distributions (like PROD at non-prime moduli) allow divide-and-conquer strategies—learning frequent outputs first, then rarer ones.
+This explains why PROD difficulty changes drastically with modulus: at primes, the output distribution is nearly uniform; at non-primes, divisibility creates structure.
+The parity and modular clustering we observe in embeddings may emerge from such divide-and-conquer: e.g., in ADD20, even+even and odd+odd both yield even results, creating exploitable structure.
+Our separation score (Fig.
+5) provides quantitative evidence: tasks inducing higher separation by divisors tend to help harder tasks more.
+Beyond arithmetic.
+Our permutation group results suggest these effects extend beyond arithmetic.
+By Cayley’s theorem, every finite group is isomorphic to a permutation subgroup, so these principles may apply to a large class of algebraic tasks.
+Connections to curriculum learning.
+Joint training and curriculum learning both exploit task relationships.
+Our transfer experiments show that sequential curricula benefit from the same mechanism: structured representations learned on easy tasks transfer to hard tasks.
+This suggests a general principle: pair tasks sharing computational primitives, not merely similar difficulty.
+Algorithmic complexity.
+Minimum model size depends not just on task complexity but on the algorithmic complexity of the discovered solution—brute-force memorization requires far more resources than efficient computation.
+Joint training appears to bias models toward efficient algorithms by constraining the representation space.
+Broader applicability.
+Prior work shows training arithmetic with text improves accuracy (Lee et al., 2024), and structured patterns help downstream tasks (Yin and Yin, 2024; Zhang et al., 2024), suggesting these principles may transfer across domains.
+Limitations.
+Our synthetic tasks may not directly transfer to natural language with ambiguous outputs, though they suggest principles for curriculum design in pretraining.
+We focus on small models; effects at larger scales remain untested.
+Our PCA-based analysis captures only linear structure, and while our output distribution hypothesis provides partial predictions, a complete theory of task compatibility remains future work.
+Impact Statement
+[…]
+Appendix A Related works
+Scaling laws for language models characterize how validation loss scales with model size (number of parameters) (Hoffmann et al., 2022), compute (FLOPs) (Muennighoff et al., 2023), dataset size (number of examples) (Hestness et al., 2017), and information capacity (in bits) (Allen-Zhu and Li, 2024), offering a foundation for model design.
 However, this perspective overlooks critical aspects of learning.
@@
 Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
-In general, language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study.
+Joint training Joint training on compositional tasks allows models to acquire the foundational primitives necessary for solving complex operations, often leading to abrupt performance improvements once all constituent skills are learned (Lubana et al., 2024; Okawa et al., 2023).
+Similarly, training on multiple arithmetic tasks has been shown to improve accuracy on individual operations, highlighting the benefits of shared representations for compositional generalization (Lee et al., 2024).
+Number representation Neural networks trained on modular addition tasks develop structured embedding spaces that reflect underlying arithmetic operations.
+Mechanistic analyses have shown that small models can exhibit ordered number patterns and implement distinct algorithmic strategies, depending on initialization and hyperparameters (Zhong et al., 2023).
+Some models converge to known solutions such as the Clock algorithm, while others discover novel procedures like the Pizza algorithm, illustrating the algorithmic diversity that can emerge from fixed training data.
+Periodic structures in the embeddings can be characterized via Fourier analysis, offering additional interpretability (Nanda et al., 2023).
+These behaviors have also been linked to grokking dynamics, where models abruptly generalize after extended training, accompanied by the emergence of structured embedding patterns (Power et al., 2022; Liu et al., 2022).
+In general language models performs poorly on symbolic mathematical tasks (Frieder et al., 2023; Dziri et al., 2023; Dave et al., 2024) such as the ListOps dataset (Nangia and Bowman, 2018) used in this study.
 Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms.
 While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively.
-Furthermore, it allows us to tune the task difficulty by combining different operations.
-Here, we analyze how language models learn arithmetic operations by training small models on the ListOps dataset.
-The ListOps dataset allows us to measure model accuracy explicitly by asking the model to solve mathematical equations.
-Using this bottom-up approach, we show that once the number of trainable parameters in the model reaches a critical point, the model is able to learn the task.
-When training models on datasets based on different levels of operation difficulty, this critical point shifts according to the complexity of the task.
-Contradictorily, we find that combining operations makes the task easier for the model than individual operations, suggesting that task diversity plays a significant role.
-lubana2024percolation discuss that structure learning drives emergent abilities.
-Seems very relevant.
-Emergence of hidden capabilities park2024emergence Another grokking paper liu2022towards claims grokking is between “ comprehension” and “memorization”.
-We may have more evidence for that.
-2 Introduction
-Scaling laws for language models kaplan2020scaling characterize how performance scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design.
-Moreover, study of emergent abilities in LLM showed wei2022emergent that various capabilities emerge at different model sizes.
-Both of these directions suggest an intimate and universal relations between model size and the emergence of abilities.
-However, these views miss important details about the role of training curriculum.
-In general language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study.
-Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms.
-While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively.
-Furthermore, it allows us to tune the task difficulty by combining different operations.
-There is always more than one way to solve a task.
-Each solution might require dramatically different amount of compute and memory resources.
-If models learn to implement specific solution algorithms, variability in resource requirements should be reflected in some way in the model size or other metrics.
-Oftentimes, when humans try to solve a task, the simplest solution (meaning, easiest to find) is a brute-force one and not particularly efficient.
-Given more thought, knowledge and experience, we often manage to find more efficient solutions.
-So a key question to ask is whether the same type of algorithmic change can be observed in machine learning models and how it can be controlled.
-How do we know which solution path a model took?
-We explore this problem through the lens of basic math operations, using the setting of the ListOps dataset (nangia2018listops).
-We show strong evidence for learning curriculum dramatically changing the algorithms learned by the models.
-Models of the same size may follow different learning paths depending on the training setup, suggesting that the tasks themselves influence the internal algorithms that emerge during training.
-It is worth noting that our observations point to more than just task complexity being the driving force.
-Complexity of a task may be defined using Kolmogorov complexity (KC) kolmogorov1965three; li2008introduction, defined as the length of the shortest code producing a desired output.
-What we show is that models may require vastly different minimum sizes to learn the same task, depending on the training curriculum.
-Additionally, we show strong evidence of a clear distinction between the internal algorithm learned by the models using different curricula.
-Such a distinct algorithm could represent learning different algorithms to solve the same problem, with each algorithm having a different resource requirement, e.g.
-brute-force versus an efficient code.
-Hence we argue the model size depends more on the algorithmic complexity, or description length of a specific algorithm, rather than the KC of the task.
-This also raises an important question: what training regimes are most effective at inducing the learning of more efficient algorithms?
-Using the ListOps dataset as our experimental framework, we investigate how small transformer models learn nested mathematical operations—specifically maximum (MAX), minimum (MIN), median (MED), and sum modulo $n$ (SUM).
-We train the models on a single or a mixture of these operations.
-We make the following observations:
-  1. 1.
-    A clear hierarchy in the model size vs learning transition: MAX and MIN are easiest to learn.
-Introducing MED and all combinations of MAX, MIN and MED require significantly more parameters to learn (2$\sim$ 2$\times$ 2 × for $n=26$).
-Next, adding SUM (mod $n$) to the mix makes it again significantly harder (1.5$\sim$ 1.5$\times$ 1.5 × for $n=26$).
-  2. 2.
-    SUM is the hardest: Paradoxically, the clear outlier as the hardest task to learn was SUM alone, not the more complex mixture of SUM and other operations.
-  3. 3.
-    Strong evidence for algorithmic diversity: The embedding layer of the pure SUM models, suggests models not learning the properties of numbers, whereas, mixed SUM models show strong number features including ordering and parity, with parity only observed in mixed SUM tasks.
-  4. 4.
-    Pretraining with MAX+MED makes SUM easier: Switching from MAX+MED to SUM has the same effect as mixed task, with much smaller models being able to learn SUM.
-We hypothesize that compared to the pure SUM task, the mixed SUM training makes it easier for the models to discover number properties of the symbols.
-This seems to lead the mixed SUM models learning an entirely different algorithm than the pure SUM model.
-It may be that the pure SUM models are memorizing the table of additions.
-Indeed, we observe a similar difficulty to learning pure SUM for models learning a scrambled SUM dataset, where we create a scrambled sum table with the right hand side of $A+B=C$ being shuffled.
-In the scrambled SUM task, any number relation between the symbols has been erased and the easiest way to learn it should be to memorize the sum table, rather than overfit the random patterns in the sum table.
-These results provide compelling evidence that joint training guides models towards finding alternative, more efficient solutions.
-They suggest there is more nuance to the scaling laws and they can be significantly affected by the training curriculum and strategies.
-3 Methodology
-We train small-scale transformer models and analyze both their performance and internal representations.
-This section outlines our methodology, detailing the dataset, model architecture, and evaluation protocols.
-Choice of task
-ListOps consists of nested mathematical equations involving operations.
-Our setup uses MAX, MIN, MED (median), and SUM (sum modulo $n$) applied to numbers (0-$n-1$).
-We conduct experiments using both $n=10$ and $n=26$, in which case we use capital letters A…Z to represent $0\dots 25$ (Appendix D.1).
-Choosing ListOps was motivated by several key factors that align with our research objectives:
-  1. 1.
-    Procedural Generation: ListOps is a synthetically generated dataset, allowing us to create a large volume of examples with controlled difficulty.
-  2. 2.
-    Exact Evaluation: The mathematical nature of ListOps operations ensures unambiguous, exact evaluation of model outputs.
-  3. 3.
-    Adjustable Task Complexity: ListOps offers a framework to modulate problem complexity by combining different operations and adjusting nesting depths.
-  4. 4.
-    Inter-task Relationships: The dataset’s multiple operations provide an opportunity to explore task synergies and interference.
-Dataset Description
-[…]
-Tokenization and CoT:
-We employ a character-based tokenization strategy for processing ListOps expressions.
-(Appendix D.2).
-We find that directly solving nested ListOps in one step can be quite challenging for transformer model (Fig.
-22) Even with a maximum of three nesting levels with three operands (inputs) we find that GPT models with over 10 million parameters still fail to learn the task.
-To enhance model performance, particularly on more complex operations like sum modulo 10, we introduced a chain of thought (CoT) approach in our training data (Appendix D.3): s(%12(%34))>(%127)>0=0e, where the ‘>’ token means one step of CoT, wherein we solve the right-most, inner-most parenthesis.
-This is similar to the scratchpad used in lee2024teaching, among others.
-The tasks become increasingly challenging to learn with increasing nesting, number of operands, and length in a predictable manner (Appendix Fig.
-27).
-Train Test Split:
-The CoT could result in many patterns appearing in one equation.
-To ensure that all the patterns in the test set are not also present in the training data, we select 100 triplets of numbers, e.g.
-$749$, to make an exclusion set.
-The training data is chosen from equations which never encounter the triplets in the exclusion set, and the test set is chosen such that each equation contains at least one excluded triplet.
-Performance Evaluation:
-We randomly select 1000 equations from the held-out test set.
-The model is prompted to generate solutions character-by-character, starting from the equation prompt (e.g., s(+3(-749))>).
-We use the output to produce two metrics: Loss: using cross-entropy computed for every character of the output.
-Accuracy: based on the number of correct answers evaluated using only the final answer, which we define as the first character after the first ‘$=$’ symbol.
-Model Architecture
-For our experiments, we employed a series of tiny GPT models, inspired by the nanoGPT architecture (karpathy2022nanogpt).
-Each model in our study uses a single attention head.
-We set the feedforward hidden dimension to four times the embedding dimension, a common practice in transformer architectures, providing sufficient complexity in the feedforward networks while keeping the model size constrained.
-By varying the embedding size and the number of transformer layers, we can create a range of models with different parameter counts.
-Figure 1: Emergence of abilities in ListOps: Each plot shows the same group of small transformer models trained on a different mix of the four operations MAX, MIN, MED, and SUM.
-Red dots are models reaching more than 50% accuracy, and blue dots are less than 50%.
-The dashed green line is a logistic fit, and the yellow star indicates the transition point at 50%.
-The x-axis is the model size (number of parameters), and the plots are sorted in ascending order of transition points.
-The bottom panel shows a bar plot of the model sizes at the transition points, with each group distinguished by a different color.
-4 Results
-To understand how language models acquire mathematical abilities, we focus on accuracy as the primary metric for observing the emergence of learning, following the approach of the emergent abilities literature as in wei2022emergent.
-While some subsequent studies have questioned the concept of emergence by examining other metrics (schaeffer2024emergent), we argue that this critique overlooks similar patterns observed in physics: In a phase transition, only some quantities may exhibit discontinuous changes.
-A metric that can capture the emergence of the new phase is called the “order parameter”.
-Hence, we choose to use accuracy as our order parameter.
-We run experiments with embedding dimension $n_{embed}$ between 4–128, and number of layers 1–4.
-In most settings, the single-layer transformers never reached 100% accuracy, so we decided to omit them.
-Similar to the emergence literature, we observe the total number of parameters to be the strongest indicator of accuracy, rather than $n_{embed}$ or depth.
-We mainly show results for modulo 26 experiments, as the patterns are more conclusive, but we also observed similar patterns for mod 10 (Appendix A.2).
-Transition to Learning.
-Figure 1 shows the transition to learning for all the 15 combinations of the four operations.
-The bottom plot shows the transition points, found using a logistic fit.
-The tasks involving MED are colored in green, and tasks involving SUM are colored in red, except for the pure SUM task, which is the curious outlier.
-We observe a clear hierarchy of tasks: adding MED to the mix makes the task significantly harder, while tasks containing SUM are even harder than MED.
-However, surprisingly, the hardest task by far is the pure SUM, requiring at least twice as many parameters as even the mixed SUM tasks.
-This is peculiar, as one might guess the complexity of the task would depend on the number of tasks mixed together and the hardness of each component task.
-Again, counter this argument, for mixed SUM tasks we observe the exact opposite: the mixture of all four tasks is easier than mixture of three, which is easier than two.
-So, something must be happening when we train multiple math tasks jointly, which makes it easier for models to learn.
-Such behavior has also been reported in prior works on language models lee2024teaching where task diversity is observed to make models smarter.
-We want to get to the bottom of how this mixture is leading to a dramatic reduction in the learning transition when training on mixed tasks in ListOps.
-Figure 2: PCA of embeddings: We choose all models which reached over $90$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure MAX, or all mixtures involving a given operation, e.g.
-Mixed SUM.
-Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do.
-Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
-Embedding Layer.
-To understand the difference between the hierarchy of transition points observed, we first examine the embedding layer of the models which learned the tasks (acc.
-$>90\%$) (see Appendices A.1 and A.2 for more).
-While the $n_{vocab}\times d_{embed}$ embedding layers of models of different sizes cannot be directly compared, we can compare them using $n_{vocab}$ dimensional principal component analysis (PCA).
-Moreover, while the PCs of individual models may be noisy, we can find aggregate PCs by first computing the $n_{vocab}\times n_{vocab}$ correlation matrices of a group of models and averaging them.
-This approach not only reduces the noise but also allows us to combine models of different sizes.
-Figure 3: PC5 and 6 of Average Mixed SUM embedding correlations perfectly separate odd and even numbers.
-An interesting picture emerges when we average models which learned each individual operation.
-Figure 2 shows the correlation matrix and top PCs for the embeddings of models trained on the groups we identified in the transition plot: 1) Pure MIN and MAX, 2) Pure and Mixed MED excluding mixture with SUM, 3) Mixed SUM, and 4) Pure SUM.
-We note that PCs for group 1 (pure MAX or MIN), as well as the mixture groups 2 and 3, exhibit clear, regular patterns.
-In contrast, the PCs for group 4 (pure SUM) look random.
-We also observe that the regular patterns are not all the same.
-For group 1, 2 and 3, the top three PCs are almost identical, with PC1 always sorting numbers, and PC2 and PC3 being the two lowest frequency waves (1 and 2 peaks).
-These three are all top eigenvector of the Laplacian of a weighted line graph.
-Interestingly, PC4 and PC5 differ among these models.
-In groups 2 and 3, MED and Mixed SUM, PC4 seems close to the next frequency (3 peaks).
-For PC5, in group 1 it looks random, in 2 it looks like the next wave’s frequency (4 peaks), but for group 3, it does something striking: it identifies parity by separating odd from even numbers.
-More specifically, a linear combination of PC5 and PC6 does parity (Fig.
-3).
-The presence of the parity mode in the embedding turns out to be a very robust feature of models trained on Mixed SUM tasks.
-We observe it not only in this grouping, but even very prominently in individual models trained on MAX+MED+SUM (Fig.
-4) and mod 10 (Fig.
-13, 20).
-Based on the observations, we pose the following hypotheses:
-  1. 1.
-    Sorting-based MAX, MIN, MED: PCA in these models show strong features of a line graph where consecutive numbers have edges and have an ordering.
-  2. 2.
-    Lookup-based Pure SUM: The lack of regularity in PCs suggest a lack of understanding of number properties, hence the algorithm might be lookup and memory-based.
-  3. 3.
-    Parity and sorting-based solution to Mixed SUM: Parity appears to play a very prominent role in the Mixed SUM algorithm learned.
-Considering that much smaller models learned Mixed SUM compared to Pure SUM, we hypothesize that models in these two groups learned entirely different algorithms for SUM.
-Mixed training seems to lead the model to develop some understanding of number properties, allowing it to use efficient and compact algorithms to solve SUM.
-Pure SUM models, in contrast, seem to be treating numbers like random symbols.
-If this is the case, we believe the models need to memorize a sum table, storing $(A,B)\to C$ for $A+B=C$.
-One way to test the hypothesis of Pure SUM finding symbolic solutions is to design a similar problem where combination of symbols $(A,B)$ turns into $C$, analogous to a shuffled sum table.
-Shuffled SUM.
-We conduct experiments with a shuffled, symmetric SUM table (i.e.
-$A+B=B+A$) in Mod 10, one set with Pure SUM and one on MAX+MED+Shuffled SUM.
-We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version (Fig.
-15).
-Additionally, we observe that MAX+MED+Shuffled SUM is again more difficult than all operations except pure sum, suggesting that the number properties played an important role in the other tasks becoming easier in mixed training.
-The PCA lacks most of the features we observed in regular numbers, save for partial ordering, as MAX and MED still use the normal number ordering (Fig.
-16).
-Coloring the numbers based on parity, we don’t see the clear parity separation we observe in the regular mod 10 experiments (Fig.
-11), as expected.
-Figure 4: Evolution of training loss, accuracy, and PCs of the embedding layer, Mod 26: The top shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 96 and 3 layers, trained either on pure SUM (blue) or on mixed MAX+MED+SUM (red).
-Models trained on MAX+MED+SUM were evaluated separately on pure MAX, MED, and SUM subsets; the corresponding accuracy curves are shown in green, purple, and orange, respectively.
-Curves are the mean on 3 runs, with shaded ±$\pm\sigma$.
-All models were trained for 20000 iterations.
-The figures beneath the main plot display PCA embeddings revealing that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss.
-The also show a prominent parity separation emerging in PC2 and PC4.
-Parity is colored by red and blue.
-In contrast, models trained solely on SUM exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
-Learning dynamics Pure SUM vs MAX+MED+SUM.
-Another piece of evidence for the difference between SUM and the rest comes from the training dynamics.
-Figure 4 shows the loss curves and the evolution of test accuracy in a pure SUM model vs MAX+MED+SUM, for models with $n_{embed}=96$ and 3 layers, which is slightly above the pure SUM learning transition.
-The mixed model was evaluated separately for each of the MAX, MED, and SUM test sets to see when each ability emerged.
-We observed a subtle but important sequence in learning: MAX was learned fastest, followed closely by MED, and then SUM.
-Crucially, the gaps between the learning of these operations were minimal, with the model beginning to grasp SUM while still perfecting MED.
-The accuracy curves also show that joint training accelerates the learning of the SUM operation, as evidenced by the much earlier rise of the orange dashed line (joint training) compared to the blue dashed line (pure SUM training), occurring at almost half the number of steps of pure SUM.
-Joint Training May Shrink the Search Space.
-The model trained on mixed data can take a very different route and converge to a different solution.
-We see that training on MAX, MIN, or MED all lead to embeddings which exhibit number properties.
-Because of the CoT steps, the model’s loss decreases when it learns any of the operations involved in the mix.
-For instance, since learning MAX is much easier than SUM, gradient descent may first learn to solve MAX, which is strongly associated with learning a representation for the digits with the correct ordering.
-This restriction of the embedding layer can make it significantly easier for the model to learn other number properties and, possibly, find a number-based algorithm for SUM, exploiting the properties of numbers instead of memorizing symbolic patterns.
-4.1 Testing the embedding restriction hypothesis.
-Figure 5: Learning SUM by finetuning MAX+MED: We train model much smaller than the sum-only learning transition (48 embedding, 2 layers).
-By switching the training data slowly from MAX+MED to pure SUM (never showing expression mixing all three) the model is able to learn SUM (blue) in this much lower parameter regime.
-In comparison, the pure SUM models (red) did not learn at this size.
-We test the hypothesis that restricting the embedding to what is learned by operations like MAX and MED could lead to smaller models learning SUM.
-To further explore this hypothesis, we designed a transfer learning experiment: 1) We first trained a model ($n_{embed}=48$ and three layers, which on pure SUM was unable to learn) on MAX and MED operations until proficiency.
-2) We then gradually introduced SUM operations, increasing their proportion in the training data from 0% to 100% over 1000 steps, while simultaneously phasing out MAX and MED.
-Key findings from this experiment include:
-  • The model began learning SUM immediately upon its introduction, despite never seeing mixed expressions (e.g., SUM with MAX or MED).
-  • Interestingly, the model started to forget MAX and MED once these operations were no longer present in the training data.
-  • Crucially, we verified that the model retained its number-like embedding structure even after MAX and MED were completely phased out.
-Perhaps most strikingly, we found that even much smaller models (embedding size 24) could learn SUM perfectly using this hybrid approach, mirroring the efficiency of MAX+MED+SUM models.
-This resulted in a model capable of performing SUM operations that was 7x smaller than the PURE SUM model, while relying on a more sophisticated, number-like internal representation.
-Attention and feedforward layers.
-To investigate the internal dynamics of the sum-only vs MAX+MED+SUM (all3) models, we focused on the final layer of a 3-layer transformer network, featuring a single attention head and an embedding dimension of 128 (Appendix E).
-We find:
-  1. 1.
-    In the attention sublayer, the MAX+MED+SUM model shows slightly higher ratio values, suggesting that attention mechanisms play a more pronounced role when the model is trained on diverse operations.
-This could indicate that the attention sublayer is capturing more complex patterns or relationships necessary for handling multiple operations.
-  2. 2.
-    Conversely, in the feedforward sublayer, the SUM model demonstrates significantly higher ratio values.
-This suggests that when trained on SUM alone, the model relies more heavily on the feedforward network for computation.
-This could imply that the SUM operation is being implemented more directly through feedforward transformations.
-  3. 3.
-    The larger effect size in the feedforward layer (-0.34) compared to the attention layer (0.18) indicates that the difference in behavior is more pronounced in the feedforward component.
-These observations suggest a trade-off in how the network allocates its computational resources.
-The MAX+MED+SUM model appears to leverage its attention mechanism more, potentially to handle the diversity of operations it was trained on.
-In contrast, the SUM model seems to channel more of its computation through the feedforward network, possibly developing a more specialized but less flexible approach to solving the sum operation.
-Figure 6: Attention patterns and layer dynamics in SUM vs MAX+MED+SUM (all3) models.
-Each panel shows (CoT) solution to a SUM modulo 10 problem, where ’>’ indicates solution steps.
-The first row shows the input sequence, with curved lines representing attention weights from Layer 2 in a 3-layer network.
-Black lines highlight attention patterns for a specific digit (shown in orange).
-Below are shown various layer metrics including the ratio of self-attention to feedforward norms (sa/ffwd), self-attention output norms (sa_out), feedforward output norms (ffwd_out), and ratios of layer outputs to inputs (sa_out/x, ffwd/x).
-Top: Model trained only on SUM operations shows attention primarily focused on parentheses and structural elements.
-Bottom: Model trained on MAX+MED+SUM (all3) shows attention strongly connecting to digits being combined in each CoT step, suggesting direct involvement in numerical computation.
-These distinct patterns suggest fundamentally different algorithms learned by each model.
-Figure 7: Comparison of layer output/input ratio distributions for models trained on MAX+MED+SUM vs pure SUM.
-Left: Attention layer ratio histogram.
-Right: Feedforward layer ratio histogram.
-The x-axis represents the log ratio of layer output norm to input norm, while the y-axis shows the density.
-The black line represents the difference MAX+MED+SUM minus SUM distributions.
+Furthermore it allows us to tune the task difficulty by combining different operations.
+Here, we investigate how language models acquire arithmetic skills by training small models on the ListOps dataset, which enables explicit evaluation through structured mathematical expressions.
+Adopting a bottom-up approach, we find that models learn to solve these tasks once the number of trainable parameters surpasses a critical threshold.
+This threshold shifts with the difficulty of the target operation (MAX = MIN << MED << ADD), indicating a dependency on task difficulty.
+Surprisingly, joint training on multiple operations often facilitates learning, outperforming models trained on individual operations (e.g.
+MAX+MED+SUM << ADD).
+This suggests that task diversity can ease optimization by promoting shared representations.
 Supplemental Material
-Appendix A Additional plots
-A.1 Modulo 26
-We define the following token vocabulary ABCDEFGHIJKLMNOPQRSTUVWXYZse()+-/%=>, where letters are mapped to integers such that A $\rightarrow$ 0, B $\rightarrow$ 1, …, Z $\rightarrow$ 25.
-Figure 8: PCA of embeddings: We choose all models which reached over $95$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure+MAX, or all mixtures involving a given operation, e.g.
-Mixed+MAX.
-Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
-Figure 9: PCA of embeddings: We choose all models which reached over $95$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure+MAX, or all mixtures involving a given operation, e.g.
-Mixed+MAX.
-Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
-Figure 10: PCA of embeddings Odd vs Even: Same plot as above, only odd numbers colored red and even colored blue.
-Mixed SUM shows a clear odd-even separation in a few of the top PCs.
-Such a separation is not clearly observed in other cases.
-Interestingly, Pure SUM approximately separates odd-even, suggesting such separation may play a role in its algorithm.
-A.2 Modulo 10
-We conducted the same experiments also on mod 10.
-The smaller number of numbers makes definitive statements about some of the patterns more challenging.
-But all the patterns we observed in mod 26 also have parallels in mod 10, including the prominent odd-even split for operations involving SUM.
-Token vocabulary: %()+-/0123456789=>es
-Figure 11: PCA of embeddings: We choose all models which reached over $90$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure MAX, or all mixtures involving a given operation, e.g.
-Mixed SUM.
-Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do.
-Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
-Figure 12: PCA of embeddings: We choose all models which reached over $95$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure+MAX, or all mixtures involving a given operation, e.g.
-Mixed+MAX.
-Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
-Figure 13: PCA of embeddings: We choose all models which reached over $95$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure+MAX, or all mixtures involving a given operation, e.g.
-Mixed+MAX.
-Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
-Figure 14: Emergence of of abilities in ListOps, Mod 10: Each plot shows the same group of small transformer models trained on different variants of ListOps.
-Each variant uses a different mix of the four operations MAX, MIN, MED, SUM.
-Red dots are model reaching more than 50% accuracy, and blue are less than 50%.
-The dashed green line is a logistic fit and the yellow star indicates the transition point at 50%.
-The x axis is the model size and the plots are sorted in ascending order of transition points.
-The bottom is a bar plot showing the model size at the transition point.
-We observe that SUM is a clear outlier, with models requiring significantly more parameters to learn SUM.
-Surprisingly, combining SUM with other operations dramatically reduces the transition point, with model less than half the size easily reaching 100% accuracy.
+Appendix B Additional plots
+B.1 Modulo 20
+Figure 10: ADD+PROD embeddings and PCs. PCs colored based modulo 4.
+Figure 11: ADD, PROD, ADD+PROD embeddings and PCs.
+B.2 Modulo prime - PROD
+Figure 12: PCs PROD modulo prime vs not prim.
+Appendix C Permutation groups.
+Figure 13: Permutation - block diagonal with two 3x3 block matrix Red indicates a cycle sub-block, while blue indicates a non-cycle block.
+Figure 14: 6x6 block diagonal permutation operations(OP, Op_TOP, OP_Bottom) OP_Top only acts on the top block matrix and OP_Bottom only acts on the bottom block matrix, while OP is an operation that acts on both the top and bottm blocks.
+Figure 15: Permutation data structure.
+Left: Operation table (OP) - block diagonal with two 3x3 block matrix.
+Middle: Operation table (OP_BOTTOM) - block diagonal with two 3x3 block matrix.
+Right: Operation table (OP_TOP) - block diagonal with two 3x3 block matrix
+Figure 16: Permutation experiments: The correlation matrix and top 2 PCA embedding f=of permutaion matrices for high performing models in two settings: Left: trianing on the full $6\times 6$ matrix multiplication (OP).
+Right: trianing on a mixture of OP, and subtasks OP_BOTTOM (multiplying bottom $3\times 3$ block, repeating the top block of first matrix), and and OP_TOP (multiply top $3\times 3$ block, repeat bottom of first matrix).
+The mixed task model reaches $90\%$ accuracy at 1.1k parameters, whereas the single OP model requires 1.9k for the same accuracy level.
+Appendix D Shuffled SUM
+The Shuffled SUM table was constructed using the following code:
+Listing 1: Generate upper triangle and diagonal matrices
+# Generate upper triangle and diagonal matrices.
+# To ensure commutativity i+j=j+i, we will transpose the
+# the upper triangle.
+upper_triangle_matrix = {(i, j): (i + j) % MOD
+for i in range(MOD)
+for j in range(MOD) if i < j}
+# To ensure uniform distribution after shuffling the values,
+# we must shuffle the diagonal separately.
+# This is because all off-diag r.h.s. are repeated twice
+# for commutativity, but not the diagonal entries.
+diagonal_matrix = {(i, i): (2 * i) % MOD for i in range(MOD)}
+# Function to shuffle values in a dictionary
+def shuffle_dict_values(d):
+keys = list(d.keys())
+values = list(d.values())
+random.shuffle(values)
+return dict(zip(keys, values))
+# Apply shuffling
+shuffled_triangle = shuffle_dict_values(upper_triangle_matrix)
+shuffled_diagonal = shuffle_dict_values(diagonal_matrix)
+As a result, there is no consistent mapping between the original and shuffled values (e.g., the number 1 does not always map to 2), making it difficult for the model to learn a deterministic transformation.
+While the shuffled sum table remains commutative, we did not enforce associativity.
+A direct check also confirmed that it does not satisfy the associative property for most triplets.
+Figure 17: Shuffle the summation table.
+D.1 Shuffled SUM modulo 10
 Figure 18: Shuffled symmetric sum table, Mod 10: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version.
@@
 There seems to be partial ordering of the numbers, but the clear wave patterns and clear parity separation is not evident.
-Figure 17: PCA of embeddings with the same sample size (67): We choose the smallest 67 models which reached over $95$% test accuracy.
-Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
-Pure+MAX, or all mixtures involving a given operation, e.g.
-Mixed+MAX.
-We observe that the PCs get slightly distorted with the smaller sample size compared to fig.
-13, but the overall structure stays the same.
-Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
-Figure 18: Parameters of continuous logistic regression fitting accuracy to number of parameters (No Early stopping).
-Figure 19: Evolution of training loss, accuracy, and principal components of cosine similarities in the embedding layer.
-Modulo 10 ListOps.
-The top main figure shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 128 and 3 layers, trained either on SUM-only data (blue) or on mixed MAX+MED+SUM data (red).
-Curves represent the mean across three independent runs, with shaded regions indicating one standard deviation.
-All models were trained for 20000 iterations.
-The red and blue boxes beneath the main plot display the average embedding representations at different training stages (indicated by vertical gray dashed lines).
-PCA reveals that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss.
-In contrast, models trained solely on SUM data exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
-This suggests that discovering an effective algorithm for the SUM operation in isolation requires significantly more exploration during training, in contrast to the more efficient joint training setting.
-Figure 20: Evolution of training loss, accuracy, and PCs of cosine similarities in the embedding layer.
-ListOps base 10.
-The top main figure shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 128 and 3 layers, trained either on SUM-only data (blue) or on mixed MAX+MED+SUM data (red).
-Curves represent the mean across three independent runs, with shaded regions indicating one standard deviation.
-All models were trained for 20000 iterations.
-The red and blue boxes beneath the main plot display PCs revealing that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss.
-In contrast, models trained solely on SUM data exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
-Appendix B Triplet Experiments
-To increase control in our experimental setup, we train models only on triplet inputs 21. We construct 1000 unique triplet samples, splitting them into 900 for training and 100 for testing.
-For joint training on the MAX, MED, and SUM operations, we use a combined training set of 2700 triplets (900 per operation) and evaluate performance on the same 100 excluded SUM samples.
-For a fair comparison, we also create a balanced training set of 900 samples by selecting 300 examples from each of MAX, MED, and SUM.
-With early stopping, models trained on the full 2700-sample mixed dataset achieve 100% accuracy on the SUM test set, while those trained on only 900 SUM samples reach approximately 50% accuracy.
-When trained for 50k steps without early stopping, models eventually learn the task, but SUM requires significantly more parameters—consistent with earlier observations.
-These results highlight that learning the SUM operation in isolation is more challenging and slower (Fig.
-21).
-Figure 21: Training only on triplet ListOps data.
-The first row shows the same group of small transformer models trained on different variants of the triplet dataset.
-Each training set is constructed from 900 unique triplets.
-The 2700 MAX+MED+SUM dataset includes all 900 triplets, each labeled with three different operations.
-The 900 MAX+MED+SUM dataset contains 300 examples per operation, randomly sampled from the 2700 set.
-The 900 SUM dataset contains only the 900 unique SUM triplets.
-In all cases, the test set comprises the same held-out 100 SUM triplets.
-Red dots indicate models that exceed 50% test accuracy, while blue dots denote models that fall below this threshold.
-The dashed green line represents a logistic fit, and the yellow star marks the transition point at 50% accuracy.
-The x-axis shows the model size (in number of parameters), and subplots are ordered by increasing transition threshold.
-The red and blue boxes beneath the main plot display the average PCA embedding representations at selected regions; red corresponds to high-accuracy models, while blue indicates low-accuracy ones.
-Models trained on the combined MAX+MED+SUM task develop a discernible structure during training, unlike models trained solely on SUM.
-Notably, augmenting MAX and MED data with only 300 SUM samples allows the model to achieve 50% accuracy on a held-out 100 SUM sample test set, using a model 10× smaller than those trained on 900 SUM samples alone, which only reach $\approx 80\%$ accuracy.
-Appendix C Related works
-Scaling laws for language models characterize how validation loss scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design.
-However, this perspective overlooks critical aspects of learning.
-Our results show that the composition of the training data—specifically, the mix of tasks—can significantly alter what and how a model learns.
-Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
-Joint training Joint training on compositional tasks allows models to acquire the foundational primitives necessary for solving complex operations, often leading to abrupt performance improvements once all constituent skills are learned lubana2024percolation; okawa2023compositional.
-Similarly, training on multiple arithmetic tasks has been shown to improve accuracy on individual operations, highlighting the benefits of shared representations for compositional generalization lee2024teaching.
-Number representation Neural networks trained on modular addition tasks develop structured embedding spaces that reflect underlying arithmetic operations.
-Mechanistic analyses have shown that small models can exhibit ordered number patterns and implement distinct algorithmic strategies, depending on initialization and hyperparameters zhong2023clock.
-Some models converge to known solutions such as the Clock algorithm, while others discover novel procedures like the Pizza algorithm, illustrating the algorithmic diversity that can emerge from fixed training data.
-Periodic structures in the embeddings can be characterized via Fourier analysis, offering additional interpretability nanda2023progress.
-These behaviors have also been linked to grokking dynamics, where models abruptly generalize after extended training, accompanied by the emergence of structured embedding patterns power2022grokking; liu2022towards.
-In general language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study.
-Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms.
-While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively.
-Furthermore it allows us to tune the task difficulty by combining different operations.
-Here, we investigate how language models acquire arithmetic skills by training small models on the ListOps dataset, which enables explicit evaluation through structured mathematical expressions.
-Adopting a bottom-up approach, we find that models learn to solve these tasks once the number of trainable parameters surpasses a critical threshold.
-This threshold shifts with the difficulty of the target operation (MAX = MIN < MED < SUM), indicating a dependency on task difficulty.
-Surprisingly, joint training on multiple operations often facilitates learning, outperforming models trained on individual operations (e.g.
-MAX+MED+SUM < SUM).
-This suggests that task diversity can ease optimization by promoting shared representations.
-Appendix D Data and processing
-D.1 Dataset Notation
-ListOps consists of nested mathematical equations involving operations such as min, max, median, and sum modulo 10 applied to single-digit numbers (0-9).
-It uses the Polish notation: (operation, [inputs]) For example: max(3,min(7,4,9))=4.
-[…]
-[…]
-[…]
-In this notation, ’s’ denotes the start of the expression, ’e’ marks the end, and parentheses indicate nesting levels.
-D.2 Tokenization
-We employ a character-based tokenization strategy for processing ListOps expressions. This approach offers several advantages:
+D.2 Shuffled SUM modulo 26
+We make a symmetric sum table ($A+B=B+A$), with a randomly shuffled right hand side, meaning where if in the table $A+B=C$, $C$ does not the actual arithmetic sum of $A+B$ modulo 26.
+We do this to test a couple of hypotheses:
   1. 1.
-    Simplicity: Character-level tokenization eliminates the need for complex tokenization rules or a large vocabulary.
+    Hypothesis 1: The SUM trained alone is not really learning the logic arithmetic of numbers, but rather memorizing the sum table.
   2. 2.
-    Generalizability: It allows the model to potentially generalize to unseen number combinations or deeper nesting levels.
-Each character in the ListOps expression, including digits, operation symbols, and structural elements (parentheses, ’s’, ’e’), is treated as a separate token.
-This granular representation enables the model to learn the syntactic structure of the expressions alongside their semantic content.
-D.3 Chain of Thought Implementation
-We find that directly solving nested ListOps in one step can be quite challenging for transformer model (Fig.
-22) Even with a maximum of three nesting levels with three operands (inputs) we find that GPT models with over 10 million parameters still fail to learn the task.
-To enhance model performance, particularly on more complex operations like sum modulo 10, we introduced a chain of thought (CoT) approach in our training data.
-This method involves providing step-by-step solutions that resolve the deepest nesting level at each step.
-For example:
-s(%12(%34))>(%127)>0=0e
-In this CoT representation:
-  • The initial expression is s(%12(%34))
-  • The first step resolves the innermost operation: (%34) becomes ‘7‘
-  • The intermediate result is shown: s(%12(7))>(%127)
-  • The process continues until the final result is reached: s(%12(7))>(%127)>0=0e
-This CoT approach serves multiple purposes: 1.
-It guides the model through the problem-solving process, mimicking human-like reasoning.
-2.
-It provides more granular supervision, potentially aiding in learning complex operations.
-3.
-It allows us to study how models learn to break down and solve nested problems.
-Our experiments show that this CoT method significantly improves model performance, particularly for the challenging sum modulo 10 operation (Fig.
-22).
-Figure 22: no - CoT vs.
-CoT.
-Providing solutions as chain-of-thought (CoT) helps models learn the tasks.
-In almost all cases, CoT accelerates learning, enabling smaller models to succeed.
-This effect is especially strong for the sum operation, which cannot be learned without CoT.
-Appendix E Observations from the norm of Attention and Feedforward outputs
-To investigate the internal dynamics of our models, we focused on the final layer of a 3-layer transformer network, featuring a single attention head and an embedding dimension of 128.
-Our analysis centered on comparing the behavior of models trained on ”ALL3” operations (MAX+MED+SUM) versus those trained solely on the ”SUM” operation.
-We introduced a novel metric to quantify the impact of different components within the network: the ratio of output to input norms for both the self-attention (SA) and feedforward (FFN) sublayers.
-Specifically, we computed:
+    Hypothesis 2: Joint training with max and med leads to learning number properties.
+If the first hypothesis holds, the shuffled sum would also require similarly high number of parameters as the normal sum and show similarly random patterns in the embedding space.
+If the second hypothesis holds, then joint training of shuffled sum should actually have detrimental effects on learning sum because number properties don’t play a role in learning shuffled sum.
+Thus we expect the jointly trained model to struggle to learn all three operations, or require significantly higher number of parameters to master shuffled sum, compared to joint training on regular sum.
+Figure 20: Shuffled symmetric sum table, Mod 26, 20k steps: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version.
+We also observe that MAX+MED+Shuffled SUM never reaches more than 80% accuracy.
+The third scatter plot from the let shows the accuracy of the MAX+MED+Shuffled SUM model on the Shuffled SUM test set.
+We see that the accuracy is very low ($\sim 20\%$ top), showing that the mixed model never learned the shuffled SUM.
+This may suggest that MAX+MED revealed number properties, but Shuffled SUM was incompatible with those properties, leading to a model that overall cannot solve the two problems (MAX+MED and Shuffled SUM) simultaneously.
+Figure 21: Shuffled symmetric sum table, Mod 26, 40 steps: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version.
+We also observe that MAX+MED+Shuffled SUM never reaches more than 80% accuracy.
+The third scatter plot from the let shows the accuracy of the MAX+MED+Shuffled SUM model on the Shuffled SUM test set.
+We see that the accuracy is very low ($\sim 20\%$ top), showing that the mixed model never learned the shuffled SUM.
+This may suggest that MAX+MED revealed number properties, but Shuffled SUM was incompatible with those properties, leading to a model that overall cannot solve the two problems (MAX+MED and Shuffled SUM) simultaneously.
+Figure 22: PCA of embeddings for Shuffled symmetric SUM, pure vs mixed with MAX and MED, Mod 26: The numbers are colored based on parity (odd is red, even is blue).
+There seems to be partial ordering of the numbers, but the clear wave patterns are not visible.
+We do almost observe parity separation in PC3 and PC4, albeit with some noise.
+It is curious that the system still learns partial parity, but evidently this feature did not allow the system to learn the shuffled SUM perfectly.
+Appendix E Deep transformer vs. Recurrent Transformer
+Figure 23: Recurrent Transformer vs.
+Deep transformer In the Rec-GPT block, there is only one transformer block that is applied recursively n times, so each output becomes the input for the next step.
+In contrast, in a traditional GPT block, there are n transformer blocks stacked sequentially on top of each other.
+Appendix F Results using Deep Transformer based models
+F.1 Experiment settings
   1. 1.
-    Attention ratio: $r_{attn}=\frac{\|SA(LN_{1}(x))\|}{\|x\|}$
-  2. 2.
-    Feedforward ratio: $r_{ffwd}=\frac{\|FFN(LN_{2}(x_{1}))\|}{\|x_{1}\|}$
-where $LN_{1}$ and $LN_{2}$ are layer normalization operations, and $x_{1}$ is the output of the self-attention sublayer.
-These ratios provide insight into how much each component modifies its input, serving as a proxy for the component’s impact on the overall computation.
-We analyzed the distribution of these ratios across a test set consisting of sum operations for both the ’ALL3’ and ’SUM’ models.
-Kernel Density Estimation (KDE) plots were used to visualize the distributions, and we employed several statistical measures to quantify the differences.
-Attention Sublayer
-The attention sublayer showed moderate but statistically significant differences between the ’ALL3’ and ’SUM’ models:
-  • Kolmogorov-Smirnov test: statistic = 0.1592, p-value < 0.0001
-  • Jensen-Shannon divergence: 0.1591
-  • Wasserstein distance: 0.0696
-  • Effect size (Cohen’s d): 0.1761
-  • 95% CI for mean difference: (0.0466, 0.0860)
-The KDE plot revealed that the ’ALL3’ model’s attention ratio distribution was more concentrated and peaked higher than the ’SUM’ model’s distribution.
-The positive effect size and confidence interval indicate that the ’ALL3’ model generally had higher attention ratios.
-Feedforward Sublayer
-The feedforward sublayer exhibited more pronounced differences:
-  • Kolmogorov-Smirnov test: statistic = 0.2461, p-value < 0.0001
-  • Jensen-Shannon divergence: 0.1617
-  • Wasserstein distance: 0.2830
-  • Effect size (Cohen’s d): -0.3379
-  • 95% CI for mean difference: (-0.3042, -0.2226)
-The KDE plot for the feedforward ratios showed a clear shift between the two distributions.
-The ’SUM’ model’s distribution was shifted towards higher values, as confirmed by the negative effect size and confidence interval.
-Interpretation
-These results reveal distinct operational patterns between models trained on ’ALL3’ operations versus those trained solely on ’SUM’:
-  1. 1.
-    In the attention sublayer, the ’ALL3’ model shows slightly higher ratio values, suggesting that attention mechanisms play a more pronounced role when the model is trained on diverse operations.
-This could indicate that the attention sublayer is capturing more complex patterns or relationships necessary for handling multiple operations.
-  2. 2.
-    Conversely, in the feedforward sublayer, the ’SUM’ model demonstrates significantly higher ratio values.
-This suggests that when trained on ’Sum’ alone, the model relies more heavily on the feedforward network for computation.
-This could imply that the ’SUM’ operation is being implemented more directly through feedforward transformations.
-  3. 3.
-    The larger effect size in the feedforward layer (-0.3379) compared to the attention layer (0.1761) indicates that the difference in behavior is more pronounced in the feedforward component.
-These observations suggest a trade-off in how the network allocates its computational resources.
-The ’ALL3’ model appears to leverage its attention mechanism more, potentially to handle the diversity of operations it was trained on.
-In contrast, the ’SUM’ model seems to channel more of its computation through the feedforward network, possibly developing a more specialized but less flexible approach to solving the sum operation.
-This analysis provides evidence that the internal dynamics of transformer models adapt significantly based on the diversity of tasks they are trained on, even when evaluated on the same type of operation (’SUM’).
-It highlights the importance of considering task diversity in understanding and optimizing neural network architectures.
-Figure 23: Comparison of layer output/input ratio distributions for models trained on all three operations (all3 - MAX+MED+SUM) versus sum operation alone (SUM).
-Left: Attention layer ratio histogram.
-Right: Feedforward layer ratio histogram.
-The x-axis represents the log ratio of layer output norm to input norm, while the y-axis shows the density.
-The black line represents the difference between the all3 and sum distributions (all3 - sum).
-These plots illustrate distinct operational patterns between the two models, with the attention layer showing increased activity in the all3 model and the feedforward layer demonstrating higher ratios in the sum model.
-Appendix F Ablation studies
-Figure 24: Random sum table.
-Joint training helps in learning even randomized sum tables.
-The numbers in the sum table follow a uniform distribution.
-We found that a randomized sum table with a non-uniform distribution is also easy to learn.
-However, a randomized sum table with the same uniform distribution remains as hard to learn as the original sum table.
-Early stopping criteria were applied during training in all simulations.
-Figure 25: Whittaker-Shannon number representation for cyclic SUM Modulo 10.
-Top row: Cyclic matrices generated using Whittaker–Shannon interpolation illustrate how digit embeddings can encode modular addition via matrix multiplication.
-Middle row (red box): Embeddings learned by a model trained on MAX+MED+SUM.
-Bottom row (blue box): Cyclic SUM representation constructed using Whittaker–Shannon numbers and digit ordering, showing a similar structure to the model’s embeddings—suggesting that the model discovers an emergent encoding of cyclic SUM.
-Appendix G Complexity: Number of operands and nesting level.
-In previous studies, task complexity has been characterized through various measures, including the number of bits required to memorize the task, which corresponds to the length of the expression dave2024investigating, the number of operands and nesting depth petruzzellis2024benchmarking, and the structure of the computational graph dziri2024faith, which captures the number of nestings.
-Here, we examine how the number of operands and nesting levels affect the learning ability of small language models.
-Focusing on the all3 task, which combines max, med, and sum operations, we manipulate complexity by varying the number of operands (arg = {3, 4, 5}) and nesting depth (depth = {3, 4, 5}).
-The length of equations, measured by the number of characters, depends on both the number of operands and the nesting level.
-We find that nesting has a greater impact: increasing the nesting level from 3 to 4 results in longer equations than increasing the number of operands from 3 to 5 at a fixed nesting level.
-We train the GPT model on the all3 dataset with all combinations of arg (number of operands) and depth (nesting levels), finding that the model’s performance correlates with the sum of operands and nesting levels (arg + depth).
-Notably, transition points tend to group together for configurations with the same sum (Fig.
-26, 27b).
-While Fig.
-27a demonstrates that the model requires more parameters to solve longer equations, it also indicates that arg + depth serves as a reliable predictor of the transition point.
-Figure 26: Learning MAX+MED+SUM operations with varying numbers of operands and nesting levels.
-The model requires more parameters as the number of operands and nesting levels increases.
-Higher nesting levels particularly demand larger model sizes to learn the task.
-We present the average of five simulations for each configuration and fit a sigmoid function, with the cross marking the middle value (transition point).
-The transition points reveal an interesting pattern: the sum of the number of operands and nesting levels groups together.
-For example, the transition points for arg-3,depth-4 (orange) and arg-4,depth-3 (red) are close to each other, as are those for arg-4,depth-5 (brown), and arg-5,depth-4 (grey).
-Early stopping criteria were applied during training in all simulations.
-(a) Average equation length vs. transition point.
-(b) Transition points.
-Figure 27: Transition point vs.
-Equation Lenght and Nesting Depth.
-(a) Transition point in function of the average equation length.
-(b) Heat plot of transition point in function number of operands and nesting depth.
-Appendix H Experiment settings.
-  1. 1.
-    Model: GPT model (karpathy2022nanogpt)
+    Model: GPT model ((Karpathy, 2022))
   2. 2.
@@
 The experiments are lightweight, requiring approximately 1.5GB of RAM, and can be executed efficiently on any recent laptop-class device.
-Appendix I Experiments on ListOps.
+F.2 Experiments on ListOps.
   1. 1.
@@
     Triplet: MAX+MED+SUM 2700 (2700 training data samples = 900 MAX + 900 MED + 900 SUM), MAX+MED+SUM 900 (900 training data samples = 300 MAX + 300 MED + 300 SUM), SUM (900 training data samples)
-Appendix J Shuffled SUM modulo 26
-We make a symmetric sum table ($A+B=B+A$), with a randomly shuffled right hand side, meaning where if in the table $A+B=C$, $C$ does not the actual arithmetic sum of $A+B$ modulo 26.
-We do this to test a couple of hypotheses:
-  1. 1.
-    Hypothesis 1: The SUM trained alone is not really learning the logic arithmetic of numbers, but rather memorizing the sum table.
-  2. 2.
-    Hypothesis 2: Joint training with max and med leads to learning number properties.
-If the first hypothesis holds, the shuffled sum would also require similarly high number of parameters as the normal sum and show similarly random patterns in the embedding space.
-If the second hypothesis holds, then joint training of shuffled sum should actually have detrimental effects on learning sum because number properties don’t play a role in learning shuffled sum.
-Thus we expect the jointly trained model to struggle to learn all three operations, or require significantly higher number of parameters to master shuffled sum, compared to joint training on regular sum.
-Figure 28: Shuffled symmetric sum table, Mod 26, 20k steps: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version.
-We also observe that MAX+MED+Shuffled SUM never reaches more than 80% accuracy.
-The third scatter plot from the let shows the accuracy of the MAX+MED+Shuffled SUM model on the Shuffled SUM test set.
-We see that the accuracy is very low ($\sim 20\%$ top), showing that the mixed model never learned the shuffled SUM.
-This may suggest that MAX+MED revealed number properties, but Shuffled SUM was incompatible with those properties, leading to a model that overall cannot solve the two problems (MAX+MED and Shuffled SUM) simultaneously.
-Figure 29: Shuffled symmetric sum table, Mod 26, 40 steps: We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version.
-We also observe that MAX+MED+Shuffled SUM never reaches more than 80% accuracy.
-The third scatter plot from the let shows the accuracy of the MAX+MED+Shuffled SUM model on the Shuffled SUM test set.
-We see that the accuracy is very low ($\sim 20\%$ top), showing that the mixed model never learned the shuffled SUM.
-This may suggest that MAX+MED revealed number properties, but Shuffled SUM was incompatible with those properties, leading to a model that overall cannot solve the two problems (MAX+MED and Shuffled SUM) simultaneously.
-Figure 30: PCA of embeddings for Shuffled symmetric SUM, pure vs mixed with MAX and MED, Mod 26: The numbers are colored based on parity (odd is red, even is blue).
-There seems to be partial ordering of the numbers, but the clear wave patterns are not visible.
-We do almost observe parity separation in PC3 and PC4, albeit with some noise.
-It is curious that the system still learns partial parity, but evidently this feature did not allow the system to learn the shuffled SUM perfectly.
+F.3 Data and processing
+F.3.1 Dataset Notation
+ListOps consists of nested mathematical equations involving operations such as min, max, median, and sum modulo 10 applied to single-digit numbers (0-9).
+It uses the Polish notation: (operation, [inputs]) For example: max(3,min(7,4,9))=4.
+max(3,min(7,4,9))=4 $\quad\Rightarrow\quad$ Polish: (max,3,(min,7,4,9))=4
+To disentangle any complexity arising from tokenization we further simplify these expression by representing the by symbols: ’++’ for max, ’-’ for min, ’//’ for median, and $\%$’ for sum modulo 10.
+For example:
+(max,3,(min,7,4,9))=4 $\quad\Rightarrow\quad$ Our notation: s(+3(-749))=4e
+In this notation, ’s’ denotes the start of the expression, ’e’ marks the end, and parentheses indicate nesting levels.
+F.3.2 Tokenization
+We employ word-level tokenization where each semantically meaningful unit is a single token:
+  • Operations: Each operation with its modulus is one token (e.g., add_20, max_10, prod_19).
+  • Numbers: Each number is one token, regardless of digit count (e.g., 3, 13, 104 are each single tokens).
+  • Symbols: Parentheses (, ), equals =, CoT separator >, and end token <eos> are each single tokens.
+Note that commas are not needed as token boundaries are implicit.
+For modulo $n$, the vocabulary size is approximately $n$ + (number of operations) + 5 (for symbols).
+This tokenization ensures that each number is atomic, preventing the model from learning spurious digit-level patterns.
+F.3.3 Chain of Thought Implementation
+Directly solving nested ListOps in one step is challenging for transformers (Fig.
+24).
+Even with maximum three nesting levels and three operands, GPT models with over 10 million parameters fail to learn.
+To address this, we use a chain-of-thought (CoT) approach that resolves the innermost operation at each step.
+For example, with the functional notation:
+add_10(1 2 add_10(3 4))>add_10(1 2 7)>0=0<eos>
+The CoT proceeds as:
+  • Initial expression: add_10(1 2 add_10(3 4))
+  • Step 1: Resolve innermost add_10(3 4) $\to$ 7
+  • Step 2: Resolve add_10(1 2 7) $\to$ 0
+  • Final: 0=0<eos>
+This CoT approach guides the model through step-by-step reasoning, provides granular supervision, and significantly improves performance on challenging operations (Fig. 24).
+Figure 24: no - CoT vs.
+CoT.
+Providing solutions as chain-of-thought (CoT) helps models learn the tasks.
+In almost all cases, CoT accelerates learning, enabling smaller models to succeed.
+This effect is especially strong for the sum operation, which cannot be learned without CoT.
+F.4 Modulo 26
+We define the following token vocabulary ABCDEFGHIJKLMNOPQRSTUVWXYZse()+-/%=>, where letters are mapped to integers such that A $\rightarrow$ 0, B $\rightarrow$ 1, …, Z $\rightarrow$ 25.
+Figure 25: PCA of embeddings: We choose all models which reached over 9595% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure+MAX, or all mixtures involving a given operation, e.g.
+Mixed+MAX.
+Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
+Figure 26: PCA of embeddings: We choose all models which reached over 9595% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure+MAX, or all mixtures involving a given operation, e.g.
+Mixed+MAX.
+Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
+Figure 27: PCA of embeddings Odd vs Even: Same plot as above, only odd numbers colored red and even colored blue.
+Mixed SUM shows a clear odd-even separation in a few of the top PCs.
+Such a separation is not clearly observed in other cases.
+Interestingly, Pure SUM approximately separates odd-even, suggesting such separation may play a role in its algorithm.
+Figure 28: PCA of embeddings: We choose all models which reached over 9090% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure MAX, or all mixtures involving a given operation, e.g.
+Mixed SUM.
+Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do.
+Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
+Figure 29: Emergence of abilities in ListOps: Each plot shows the same group of small transformer models trained on a different mix of the four operations MAX, MIN, MED, and SUM.
+Red dots are models reaching more than 50% accuracy, and blue dots are less than 50%.
+The dashed green line is a logistic fit, and the yellow star indicates the transition point at 50%.
+The x-axis is the model size (number of parameters), and the plots are sorted in ascending order of transition points.
+The bottom panel shows a bar plot of the model sizes at the transition points, with each group distinguished by a different color.
 Figure 30: Emergence of abilities in ListOps, 40k steps: Long training.
@@
 The bottom panel shows a bar plot of the model sizes at the transition points, with each group distinguished by a different color.
-Appendix K Attention and Feedforward outputs modulo 10
+F.5 Modulo 10
+We conducted the same experiments also on mod 10.
+The smaller number of numbers makes definitive statements about some of the patterns more challenging.
+But all the patterns we observed in mod 26 also have parallels in mod 10, including the prominent odd-even split for operations involving SUM.
+Token vocabulary: %()+-/0123456789=>es
+Figure 31: PCA of embeddings: We choose all models which reached over 9090% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure MAX, or all mixtures involving a given operation, e.g.
+Mixed SUM.
+Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do.
+Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
+Figure 32: PCA of embeddings: We choose all models which reached over 9595% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure+MAX, or all mixtures involving a given operation, e.g.
+Mixed+MAX.
+Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
+Figure 33: PCA of embeddings: We choose all models which reached over 9595% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure+MAX, or all mixtures involving a given operation, e.g.
+Mixed+MAX.
+Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
+Figure 34: Emergence of of abilities in ListOps, Mod 10: Each plot shows the same group of small transformer models trained on different variants of ListOps.
+Each variant uses a different mix of the four operations MAX, MIN, MED, SUM.
+Red dots are model reaching more than 50% accuracy, and blue are less than 50%.
+The dashed green line is a logistic fit and the yellow star indicates the transition point at 50%.
+The x axis is the model size and the plots are sorted in ascending order of transition points.
+The bottom is a bar plot showing the model size at the transition point.
+We observe that SUM is a clear outlier, with models requiring significantly more parameters to learn SUM.
+Surprisingly, combining SUM with other operations dramatically reduces the transition point, with model less than half the size easily reaching 100% accuracy.
+Figure 35: PCA of embeddings with the same sample size (67): We choose the smallest 67 models which reached over 9595% test accuracy.
+Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g.
+Pure+MAX, or all mixtures involving a given operation, e.g.
+Mixed+MAX.
+We observe that the PCs get slightly distorted with the smaller sample size compared to fig.
+33, but the overall structure stays the same.
+Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
+Figure 36: Parameters of continuous logistic regression fitting accuracy to number of parameters (No Early stopping).
+Figure 37: Evolution of training loss, accuracy, and principal components of cosine similarities in the embedding layer.
+Modulo 10 ListOps.
+The top main figure shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 128 and 3 layers, trained either on SUM-only data (blue) or on mixed MAX+MED+SUM data (red).
+Curves represent the mean across three independent runs, with shaded regions indicating one standard deviation.
+All models were trained for 20000 iterations.
+The red and blue boxes beneath the main plot display the average embedding representations at different training stages (indicated by vertical gray dashed lines).
+PCA reveals that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss.
+In contrast, models trained solely on SUM data exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
+This suggests that discovering an effective algorithm for the SUM operation in isolation requires significantly more exploration during training, in contrast to the more efficient joint training setting.
+F.6 Triplet Experiments
+To increase control in our experimental setup, we train models only on triplet inputs 38. We construct 1000 unique triplet samples, splitting them into 900 for training and 100 for testing.
+For joint training on the MAX, MED, and SUM operations, we use a combined training set of 2700 triplets (900 per operation) and evaluate performance on the same 100 excluded SUM samples.
+For a fair comparison, we also create a balanced training set of 900 samples by selecting 300 examples from each of MAX, MED, and SUM.
+With early stopping, models trained on the full 2700-sample mixed dataset achieve 100% accuracy on the SUM test set, while those trained on only 900 SUM samples reach approximately 50% accuracy.
+When trained for 50k steps without early stopping, models eventually learn the task, but SUM requires significantly more parameters—consistent with earlier observations.
+These results highlight that learning the SUM operation in isolation is more challenging and slower (Fig.
+38).
+Figure 38: Training only on triplet ListOps data.
+The first row shows the same group of small transformer models trained on different variants of the triplet dataset.
+Each training set is constructed from 900 unique triplets.
+The 2700 MAX+MED+SUM dataset includes all 900 triplets, each labeled with three different operations.
+The 900 MAX+MED+SUM dataset contains 300 examples per operation, randomly sampled from the 2700 set.
+The 900 SUM dataset contains only the 900 unique SUM triplets.
+In all cases, the test set comprises the same held-out 100 SUM triplets.
+Red dots indicate models that exceed 50% test accuracy, while blue dots denote models that fall below this threshold.
+The dashed green line represents a logistic fit, and the yellow star marks the transition point at 50% accuracy.
+The x-axis shows the model size (in number of parameters), and subplots are ordered by increasing transition threshold.
+The red and blue boxes beneath the main plot display the average PCA embedding representations at selected regions; red corresponds to high-accuracy models, while blue indicates low-accuracy ones.
+Models trained on the combined MAX+MED+SUM task develop a discernible structure during training, unlike models trained solely on SUM.
+Notably, augmenting MAX and MED data with only 300 SUM samples allows the model to achieve 50% accuracy on a held-out 100 SUM sample test set, using a model 10× smaller than those trained on 900 SUM samples alone, which only reach $\approx 80\%$ accuracy.
+Appendix G Observations from the norm of Attention and Feedforward outputs
+To investigate the internal dynamics of our models, we focused on the final layer of a 3-layer transformer network, featuring a single attention head and an embedding dimension of 128.
+Our analysis centered on comparing the behavior of models trained on ”ALL3” operations (MAX+MED+SUM) versus those trained solely on the ”SUM” operation.
+We introduced a novel metric to quantify the impact of different components within the network: the ratio of output to input norms for both the self-attention (SA) and feedforward (FFN) sublayers.
+Specifically, we computed:
+  1. 1.
+    Attention ratio: $r_{attn}=\frac{\|SA(LN_{1}(x))\|}{\|x\|}$
+  2. 2.
+    Feedforward ratio: $r_{ffwd}=\frac{\|FFN(LN_{2}(x_{1}))\|}{\|x_{1}\|}$
+where $LN_{1}$ and $LN_{2}$ are layer normalization operations, and $x_{1}$ is the output of the self-attention sublayer.
+These ratios provide insight into how much each component modifies its input, serving as a proxy for the component’s impact on the overall computation.
+We analyzed the distribution of these ratios across a test set consisting of sum operations for both the ’ALL3’ and ’SUM’ models.
+Kernel Density Estimation (KDE) plots were used to visualize the distributions, and we employed several statistical measures to quantify the differences.
+Attention Sublayer
+The attention sublayer showed moderate but statistically significant differences between the ’ALL3’ and ’SUM’ models:
+  • Kolmogorov-Smirnov test: statistic = 0.1592, p-value ¡ 0.0001
+  • Jensen-Shannon divergence: 0.1591
+  • Wasserstein distance: 0.0696
+  • Effect size (Cohen’s d): 0.1761
+  • 95% CI for mean difference: (0.0466, 0.0860)
+The KDE plot revealed that the ’ALL3’ model’s attention ratio distribution was more concentrated and peaked higher than the ’SUM’ model’s distribution.
+The positive effect size and confidence interval indicate that the ’ALL3’ model generally had higher attention ratios.
+Feedforward Sublayer
+The feedforward sublayer exhibited more pronounced differences:
+  • Kolmogorov-Smirnov test: statistic = 0.2461, p-value ¡ 0.0001
+  • Jensen-Shannon divergence: 0.1617
+  • Wasserstein distance: 0.2830
+  • Effect size (Cohen’s d): -0.3379
+  • 95% CI for mean difference: (-0.3042, -0.2226)
+The KDE plot for the feedforward ratios showed a clear shift between the two distributions.
+The ’SUM’ model’s distribution was shifted towards higher values, as confirmed by the negative effect size and confidence interval.
+Interpretation
+These results reveal distinct operational patterns between models trained on ’ALL3’ operations versus those trained solely on ’SUM’:
+  1. 1.
+    In the attention sublayer, the ’ALL3’ model shows slightly higher ratio values, suggesting that attention mechanisms play a more pronounced role when the model is trained on diverse operations.
+This could indicate that the attention sublayer is capturing more complex patterns or relationships necessary for handling multiple operations.
+  2. 2.
+    Conversely, in the feedforward sublayer, the ’SUM’ model demonstrates significantly higher ratio values.
+This suggests that when trained on ’Sum’ alone, the model relies more heavily on the feedforward network for computation.
+This could imply that the ’SUM’ operation is being implemented more directly through feedforward transformations.
+  3. 3.
+    The larger effect size in the feedforward layer (-0.3379) compared to the attention layer (0.1761) indicates that the difference in behavior is more pronounced in the feedforward component.
+These observations suggest a trade-off in how the network allocates its computational resources.
+The ’ALL3’ model appears to leverage its attention mechanism more, potentially to handle the diversity of operations it was trained on.
+In contrast, the ’SUM’ model seems to channel more of its computation through the feedforward network, possibly developing a more specialized but less flexible approach to solving the sum operation.
+This analysis provides evidence that the internal dynamics of transformer models adapt significantly based on the diversity of tasks they are trained on, even when evaluated on the same type of operation (’SUM’).
+It highlights the importance of considering task diversity in understanding and optimizing neural network architectures.
+Figure 39: Comparison of layer output/input ratio distributions for models trained on all three operations (all3 - MAX+MED+SUM) versus sum operation alone (SUM).
+Left: Attention layer ratio histogram.
+Right: Feedforward layer ratio histogram.
+The x-axis represents the log ratio of layer output norm to input norm, while the y-axis shows the density.
+The black line represents the difference between the all3 and sum distributions (all3 - sum).
+These plots illustrate distinct operational patterns between the two models, with the attention layer showing increased activity in the all3 model and the feedforward layer demonstrating higher ratios in the sum model.
+Figure 40: Attention patterns and layer dynamics in SUM vs MAX+MED+SUM (all3) models.
+Each panel shows (CoT) solution to a SUM modulo 10 problem, where ’¿’ indicates solution steps.
+The first row shows the input sequence, with curved lines representing attention weights from Layer 2 in a 3-layer network.
+Black lines highlight attention patterns for a specific digit (shown in orange).
+Below are shown various layer metrics including the ratio of self-attention to feedforward norms (sa/ffwd), self-attention output norms (sa_out), feedforward output norms (ffwd_out), and ratios of layer outputs to inputs (sa_out/x, ffwd/x).
+Top: Model trained only on SUM operations shows attention primarily focused on parentheses and structural elements.
+Bottom: Model trained on MAX+MED+SUM (all3) shows attention strongly connecting to digits being combined in each CoT step, suggesting direct involvement in numerical computation.
+These distinct patterns suggest fundamentally different algorithms learned by each model.
 Figure 41: Attention layer and Feedforward layer ratio histogram.
@@
 Each column shows the ratio histogram in different attention blocks.
-The model had 128 embedding and 3 layers and all models rechec 99% accuracy.
+The model had 128 embedding and 3 layers and all models reaches 99% accuracy.
 Figure 42: The mean attention layer ratio vs the feedforward layer ratio. Each plots show the means in different layers.
@@
 The title of the figures contain the Effective size showing the difference between the model.
+Appendix H Ablation Studies
+H.1 Number of Attention Heads
+[…]
+H.2 Deep vs. Recurrent Architecture
+Our primary architecture uses a recurrent transformer: a single transformer block applied iteratively, feeding its output back as input for the next iteration.
+We compared this to standard deep transformers with 2–4 stacked layers and found qualitatively similar results for the task difficulty hierarchy and joint training benefits (see Section F).
+The recurrent architecture was chosen for two reasons: (1) it converges faster during training, and (2) it produces more structured, interpretable embeddings that facilitate our PCA analysis.
+The key finding—that joint training with compatible tasks reduces model size requirements—holds for both architectures.
+Appendix I Additional Scaling Results
+Figure 44 shows the full scaling results across all task combinations tested, complementing the selected comparisons in the main text.
+Figure 44: Full scaling results across moduli.
+Transition points (minimum model size to learn) as a function of modulo $n$ for various task combinations.
+Prime moduli are shown in white text.
+Key observations: (1) PROD-prime is consistently hard; (2) ADD+PROD helps ADD for non-prime moduli; (3) MAX+PROD helps PROD even for primes; (4) ADD+MAX+PROD at primes can be harder than individual tasks.
+Appendix J Additional Curriculum Experiments
+Figure 45 shows the full curriculum training results including loss, accuracy, and mix ratio curves comparing baseline ADD training with the ADD+PROD $\to$ ADD curriculum.
+Figure 45: Full curriculum comparison: ADD+PROD $\to$ ADD.
+Left: Validation loss over training.
+Middle: Accuracy over training—baseline ADD (blue) stays at $\sim$15%, curriculum (orange) reaches near-perfect accuracy.
+Right: Curriculum mix ratio showing transition from ADD+PROD to pure ADD.
+In addition, we tested transfer from MAX+MED to ADD.
+MAX+MED $\to$ ADD transfer.
+We trained a small model ($n_{\text{embed}}=48$, below the pure ADD transition) on MAX+MED until proficiency, then gradually transitioned to pure ADD (never showing mixed MAX+MED+ADD expressions).
+Key findings:
+  • The model learns ADD immediately upon introduction, achieving high accuracy.
+  • The model forgets MAX and MED as they are phased out, but retains the structured embedding.
+[…]
+This confirms that the curriculum learning effect is robust across different task combinations: structured representations learned from easy tasks enable efficient learning of hard tasks.
+Figure 46: Transfer from MAX+MED to ADD.
+A model too small to learn pure ADD ($n_{\text{embed}}=48$) successfully learns ADD after pretraining on MAX+MED.
+Blue: curriculum (MAX+MED $\to$ ADD) reaches high accuracy.
+Red: baseline (pure ADD) fails to learn.
+Appendix K Complexity: Number of operands and nesting level.
+In previous studies, task complexity has been characterized through various measures, including the number of bits required to memorize the task, which corresponds to the length of the expression (Dave et al., 2024), the number of operands and nesting depth (Petruzzellis et al., 2024), and the structure of the computational graph (Dziri et al., 2023), which captures the number of nestings.
+Here, we examine how the number of operands and nesting levels affect the learning ability of small language models.
+Focusing on the all3 task, which combines max, med, and sum operations, we manipulate complexity by varying the number of operands (arg = {3, 4, 5}) and nesting depth (depth = {3, 4, 5}).
+The length of equations, measured by the number of characters, depends on both the number of operands and the nesting level.
+We find that nesting has a greater impact: increasing the nesting level from 3 to 4 results in longer equations than increasing the number of operands from 3 to 5 at a fixed nesting level.
+We train the GPT model on the all3 dataset with all combinations of arg (number of operands) and depth (nesting levels), finding that the model’s performance correlates with the sum of operands and nesting levels (arg + depth).
+Notably, transition points tend to group together for configurations with the same sum (Fig.
+47, 48b).
+While Fig.
+48a demonstrates that the model requires more parameters to solve longer equations, it also indicates that arg + depth serves as a reliable predictor of the transition point.
+Figure 47: Learning MAX+MED+SUM operations with varying numbers of operands and nesting levels.
+The model requires more parameters as the number of operands and nesting levels increases.
+Higher nesting levels particularly demand larger model sizes to learn the task.
+We present the average of five simulations for each configuration and fit a sigmoid function, with the cross marking the middle value (transition point).
+The transition points reveal an interesting pattern: the sum of the number of operands and nesting levels groups together.
+For example, the transition points for arg-3,depth-4 (orange) and arg-4,depth-3 (red) are close to each other, as are those for arg-4,depth-5 (brown), and arg-5,depth-4 (grey).
+Early stopping criteria were applied during training in all simulations.
+(a) Average equation length vs. transition point.
+(b) Transition points.
+Figure 48: Transition point vs.
+Equation Lenght and Nesting Depth.
+(a) Transition point in function of the average equation length.
+(b) Heat plot of transition point in function number of operands and nesting depth.
+Appendix L Clustering in the embedding space
+Figure 49: ADD, PROD and ADD+PRD in MOD 42 PC embeddings
+Figure 50: ADD, PROD and ADD+PRD in MOD 43 PC embeddings
+Figure 51: ADD, PROD and ADD+PRD in MOD 45 PC embeddings
+Figure 52: Clustering in the PC space.
+We measure how well modulo-2, 3, 4, 5, and 6 divisor clusters separate in the PC space for the ADD, PROD, and ADD+PROD models on MOD 42, 43, and 45 (for runs with accuracy $80\%$).
+Cluster quality is computed from the mean position and standard deviation of each modulo group, and the pairwise distances between them.
+For MOD 42, strong clustering appears for mod-2, 3, 6, and sometimes 4—precisely the divisors of 42.
+Strong PROD signals almost always produce strong ADD+PROD signals, whereas strong ADD signals without PROD support do not transfer.
+For MOD 43, no model shows strong clustering; both ADD and PROD struggle and do not help each other.
+For MOD 45, strong clustering occurs for mod-3 and mod-5, and occasionally mod-6 (via divisor 3).
+We also observe that the mixed model only exhibits strong clustering when PROD does as well, indicating that the PROD embedding drives the shared structure and helps ADD.
+Figure 53: Clustering in the PC space by averaging over the first 5 PCs.
+We measure how well modulo-2, 3, 4, 5, and 6 divisor clusters separate in the PC space for the ADD, PROD, and ADD+PROD models on MOD 42, 43, and 45 (for runs with accuracy $80\%$).
+Cluster quality is computed from the mean position and standard deviation of each modulo group, and the pairwise distances between them, averaged over the first five PCs.
+For MOD 42, strong clustering appears for mod-2, 3, and 4 (could be an artifact of 2) — precisely the divisors of 42.
+For MOD 43, we find no strong clustering compared to MOD 42 and MOD 45.
+For MOD 45, strong clustering occurs for mod-3 and mod-5, as expected from the divisors of 45.
+We also observe that the mixed model exhibits strong clustering only when PROD does as well, indicating that the PROD embedding drives the shared structure and supports ADD.